### v1

# 3M-Diffusion: Latent Multi-Modal Diffusion for Text-Guided Generation of Molecular Graphs

Abstract
Generating molecules with desired properties is a critical task with broad applications in drug discovery and materials design. Inspired by recent advances in large language models, there is a growing interest in using natural language descriptions of molecules to generate molecules with the desired properties. Most existing methods focus on generating molecules that precisely match the text description. However, practical applications call for methods that generate diverse, and ideally novel, molecules with the desired properties. We propose 3M-Diffusion, a novel multi-modal molecular graph generation method, to address this challenge. 3M-Diffusion first encodes molecular graphs into a graph latent space aligned with text descriptions. It then reconstructs the molecular structure and atomic attributes based on the given text descriptions using the molecule decoder. It then learns a probabilistic mapping from the text space to the latent molecular graph space using a diffusion model. The results of our extensive experiments on several datasets demonstrate that 3M-Diffusion can generate high-quality, novel and diverse molecular graphs that semantically match the textual description provided. Code is released at https://github.com/huaishengzhu/3MDiffusion.
Machine Learning, ICML
1 Introduction
Generating molecules with the desired properties is a critical task with broad applications in drug discovery and materials design (Hajduk & Greer, 2007; Mandal et al., 2009; Pyzer-Knapp et al., 2015). There is a growing interest in generative models to automate this task (You et al., 2018; Jin et al., 2018, 2020; Bjerrum & Threlfall, 2017).  Kusner et al. (2017) proposed a variational autoencoder that uses parse trees for probabilistic context free grammars to encode discrete sequences that specify the Simplified Molecular-Input Line-entry System (SMILES) (Weininger, 1988) notation of molecules use the resulting parse trees to generate SMILES codes that describe molecular graphs. Edwards et al. (2021) formulated molecule retrieval from textual description as a cross-lingual retrieval task. Recent advances in language models have led to innovative approaches to generate molecules from human-friendly textual descriptions of their properties, substructures, and biochemical activity (Edwards et al., 2022; Zeng et al., 2022; Zhao et al., 2023b; Fang et al., 2023a). These approaches learn mappings from textual descriptions to chemical structures using single transformer-based cross-lingual language model for molecule generation. Specifically, with the given text description of the molecule, these approaches produce SMILES strings describing the molecular graph.
Figure 1: The overview of 3M-diffusion, with a molecule graph encoder/decoder and a latent diffusion model conditioned on a prior (an aligned text encoder). The details of alignment in the text/graph encoders and diffusion model are given in Section 4.
Figure 2: Qualitative comparisons to the MolT5-large in terms of generated molecules on ChEBI-20. Compared with the SOTA method MolT5-large, our generated results are more diverse and novel with maintained semantics in textual prompt.
Despite promising initial results, such approaches suffer from some important limitations: (i) SMILES encoding is degenerate in that the a given molecular graph can have multiple SMILE encodings. For example, the structure of ethanol can be specified using CCO, OCC or C(O)C. Although a unique canonical SMILES representation can be produced for a given graph, the procedure used to do so relies on several arbitrary choices. Not surprisingly, molecules with identical chemical graphs may map to significantly distinct SMILES strings, making SMILES notation less than ideal for generating molecules from their textual descriptions (Jin et al., 2018). Important chemical properties corresponding to the substructures of molecules are more straightforward to express using molecular graphs rather than linear SMILES representations (Jin et al., 2018; Du et al., 2022); (ii) Using language models to generate diverse high-quality molecules matching a description is challenging, because the decoding process used such as beam search or greedy search often produces candidates that are often too similar to each other as shown in Figure 2. Hence, there is an urgent need for better approaches to generating diverse, novel molecular graphs from a given textual description  (Brown et al., 2019). Against this background, this paper aims to answer the following research question: How can we generate a diverse and novel collection of molecular graphs that match a given textual description?
In this work, we introduce 3M-Diffusion, a novel multi-modal molecular diffusion method, for generating high-quality, diverse, novel molecular graphs given a textual description. Inspried by text-guided image generation using a diffusion model on the latent space (Rombach et al., 2022), 3M-Diffusion is trained to generate molecules from the latent distribution of a molecular graph autoencoder (See Figure 1). This enables the diffusion process to focus on the high-level semantics of molecules. We find that this approach is particularly well-suited for discrete molecular graph modality because it offloads the challenge of modeling a discrete and diverse distribution to the autoencoder and simplifies the diffusion process by restricting it to the continuous, latent feature space. Since graph data consists of both node attributes and graph structure, which differs significantly from the sequential and contextual nature of text data, to align the latent space of the text and molecular graph, 3M-diffusion first pretrains molecular graph and text encoders through contrastive learning on a large dataset of molecule-text pairs, enabling generate high-quality molecular graphs that match the given textual descriptions. Our extensive experiments demonstrate that 3M-Diffusion can generate high-quality, novel, and diverse molecular graphs that semantically match the textual description provided.
Key Contributions. The key contributions of this paper are as follows: (i) To the best of our knowledge, 3M-Diffusion offers the first multimodal diffusion approach to generating molecular graphs from their textual descriptions, surpassing the limitations of state-of-the-art (SOTA) methods for this problem. (ii) 3M-Diffusion aligns the latent spaces of molecular graphs and textual descriptions to offer a text-molecule aligned latent diffusion model to generate higher-quality, diverse, and novel molecular graphs that match the textual description provided; and (iii) We present results of extensive experiments using four real-world text-based molecular graphs generation benchmarks that show that 3M-Diffusion outperforms SOTA methods for both text-guided and unconditional molecular graph generation. In particular, 3M-Diffusion achieves 146.27% novelty, 130.04% diversity relative improvement over the SOTA method with maintained semantics in textual prompt on the PCDes dataset.
2 Related Work
2.1 Molecule Generation
Existing molecular graph generation methods can be broadly grouped into two categories: text-based and graph-based models. Text-based models typically utilize the SMILES (Weininger, 1988) or SELFIES (Krenn et al., 2022) linear strings to describe each molecule  (Bjerrum & Threlfall, 2017; Gómez-Bombarelli et al., 2018; Kusner et al., 2017; Flam-Shepherd et al., 2022; Fang et al., 2023b; Grisoni, 2023). Because a given molecular graph can have several linear string representations and slight changes to the string representation can yield significant changes in molecular graph being described, such linear string encodings are far from ideal for learning generative models for producing molecular graph from textual descriptions, several authors have used deep generative models, including graph variational autoencoders (VAEs) (You et al., 2018; Jin et al., 2018; Xiao et al., 2021; Jin et al., 2020; Kong et al., 2022; Liu et al., 2018; Diamant et al., 2023), normalizing flows (Madhawa et al., 2019; Zang & Wang, 2020; Luo et al., 2021), generative adversarial networks (Guarino et al., 2017; Maziarka et al., 2020), diffusion models (Niu et al., 2020; Vignac et al., 2022; Jo et al., 2022) and autoregressive models (You et al., 2018; Xiao & Wang, 2021; Popova et al., 2019; Goyal et al., 2020) to learn distributions of molecular graphs from large databases of known graphs. However, many of these studies focus on generating molecular graphs with specified low-level properties, such as logP (the octanol-water partition coefficient) as opposed to those that match high-level textual descriptions of molecules (Edwards et al., 2022). However, practical applications, such as drug design, call for effective approaches to generating diverse and novel molecular graphs that match a given textual description.
2.2 Text-guided Generation of Molecular Graphs
Advances in deep learning and language models have inspired a growing body of work on applications of such models to problems in the molecular sciences  (Schwaller et al., 2019; Xiao et al., 2023; Toniato et al., 2021; Xiao et al., 2024; Zhao et al., 2023a; Seidl et al., 2023; Schwaller et al., 2021; Vaucher et al., 2020). Advances in text-guided generation of images, videos, and audios  (Radford et al., 2021; Rombach et al., 2022; OpenAI, 2023; Yin et al., 2023), together with the flexibility offered by natural language descriptions of molecular graphs have inspired a growing body of work on text-guided generation of molecular graphs (Christofidellis et al., 2023; Edwards et al., 2022; Fang et al., 2023a; Zeng et al., 2022; Su et al., 2022). For example, MolT5 (Edwards et al., 2021, 2022) and ChemT5 (Christofidellis et al., 2023) pre-train on a substantial volume of unlabeled language text and SMILES strings for tasks such as molecule captioning and text-based molecule generation (Christofidellis et al., 2023; Edwards et al., 2022). However, as noted above, the degenerate nature of linear SMILES string representation of molecular graphs makes it less ideal for generating molecules from their textual descriptions. To address this problem, we propose 3M-Diffusion, aligning latent representations of textual descriptions for molecules with the latent representations of the graphs themselves, offering a powerful approach for text-guided molecule generation.
3 Preliminaries
3.1 Problem Definition
In this paper, we consider generative modeling of 2D molecular graphs from scratch. Each molecule is represented as a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}=\left\{v_{1},\ldots,v_{|\mathcal{V}|}\right\}$ is the set of $|\mathcal{V}|$ nodes and $\mathcal{E}$ is the set of edges. Let $\mathbf{X}=\left[\mathbf{x}_{1},\mathbf{x}_{2},\cdots,\mathbf{x}_{|\mathcal{V}\mid}\right]\in\mathbb{R}^{\left[\mathcal{V}\mid\times D_{x}\right.}$ be the node attribute matrix, where $x_{i}$ is the $D_{x}$-dimensional one-hot encoding feature vector of $v_{i}$, such as atomic type and chirality type. Similarly, a tensor $\mathbf{A}\in\mathbb{R}^{|\mathcal{V}|\times|\mathcal{V}|\times b}$ groups the one-hot encoding $\mathbf{e}_{ij}$ of each edge, with each entry representing a distinct edge type (bond types for molecule), while treating the absence of an edge as a specific edge type. Given the dataset $\mathcal{D}=\{(\mathcal{G}_{i},\mathcal{T}_{i})_{i=1}^{N}\}$ where molecule $\mathcal{G}$ has specific text descriptions $\mathcal{T}$ characterizing the features of the molecule, our goal is to learn conditional generation models $p_{\theta}(\mathcal{G}\mid\mathcal{T})$ for generating novel, diverse, and valid molecules while simultaneously aligning with the desired features outlined in text descriptions $\mathcal{T}$. For brevity, in what follows, we omit the input subscript $i$ unless otherwise stated.
3.2 Diffusion Models
Diffusion models (DMs) (Ho et al., 2020; Song et al., 2020) are latent variable models that represent the data $z_{0}$ as Markov chains $\mathbf{z}_{T}\cdots\mathbf{z}_{0}$, where intermediate variables share the same dimension. DMs involve a forward diffusion process $q\left(\mathbf{z}_{1:T}\mid\mathbf{z}_{0}\right)=\prod_{t=1}^{T}q\left(\mathbf{z}_{t}\mid\mathbf{z}_{t-1}\right)$ that systematically adds Gaussian noise to samples and a reverse process $p_{\theta}\left(\mathbf{z}_{0:T}\right)=p\left(\mathbf{z}_{T}\right)\prod_{t=1}^{T}p_{\theta}\left(\mathbf{z}_{t-1}\mid\mathbf{z}_{t}\right)$ that iteratively ”denoises” samples from the Gaussian distribution to generate samples from the data distribution. A formal description of diffusion models is given in the Appendix A. To ensure training stability, we can use a simple regression objective with a denoising network $\hat{\mathbf{z}}_{\theta}$ with input $(\mathbf{x}_{t},t)$:
$\mathcal{L}_{\text{diff}}=\mathbb{E}_{t,\mathbf{x}_{0},\epsilon}\left[\lambda_{t}\left\|\hat{\mathbf{z}}_{\theta}\left(\sqrt{\alpha_{t}}\mathbf{z}_{0}+\sqrt{1-\alpha_{t}}{\bm{\epsilon}},t\right)-\mathbf{z}_{0}\right\|_{2}^{2}\right],$ (1)
where $t$ is the timestep, $\alpha_{t}\in[0,1]$ is the noise schedule and $\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\bm{I})$ is Gaussian noise and $\lambda_{t}$ is a time-dependent weighting term. Intuitively, the denoising network is trained to denoise a noisy state, $\mathbf{z}_{t}=\sqrt{\alpha_{t}}\mathbf{z}_{0}+\sqrt{1-\alpha_{t}}{\bm{\epsilon}}$, aiming to reconstruct the clean data $\mathbf{z}_{0}$ with Equation (1) that prioritizes specific times $t$. After training, we can draw samples with $\hat{\mathbf{z}}_{\theta}$ by the iterative ancestral sampling (Ho et al., 2020):
$\mathbf{z}_{t-1}$ $=\tfrac{\sqrt{\alpha_{t-1}}\left(1-\alpha_{t\mid t-1}\right)}{1-\alpha_{t}}\hat{\mathbf{z}}_{\theta}(\mathbf{x}_{t},t)$ (2)
$+\tfrac{\sqrt{\alpha_{t\mid t-1}}\left(1-\alpha_{t-1}\right)}{1-\alpha_{t}}\mathbf{z}_{t}+\sigma(t-1,t){\bm{\epsilon}},$
where $\alpha_{t\mid t-1}$ represents $\alpha_{t}/\alpha_{t-1}$ and $\sigma(t-1,t)=(1-\alpha_{t-1})\left(1-\alpha_{t\mid t-1}\right)/(1-\alpha_{t})$. The sampling chain is initialized from Gaussian prior $\mathbf{z}_{T}\sim p\left(\mathbf{z}_{T}\right)=\mathcal{N}\left(\mathbf{z}_{T};\mathbf{0},\bm{I}\right)$.
4 Latent Multi-Modal Diffusion for Molecules
Our 3M-Diffusion in Figure 1 aims to learn a probabilistic mapping from the latent space of textual descriptions to a latent space of molecular graphs for text-guided generation of molecular graphs. However, direct mapping between these two latent representations using a diffusion model faces a key hurdle, namely, large mismatch between the latent spaces of textual descriptions and of molecular graphs. To overcome this hurdle, 3M-Diffusion adopts a two-part approach using a text-molecule aligned variational autoencoder and a multi-modal molecule latent diffusion. The former learns to align the representation of molecular graphs with that of their textual descriptions using contrastive learning, followed by a decoder that maps the latent representation back to the corresponding molecular graphs. The latter works with the text-aligned latent representation of molecular graphs to learn a conditional generative model, that maps textual descriptions to latent representations of molecular graphs. As we will see later, 3M-Diffusion is able to produce a diverse collection of novel molecular graphs that match the given textual descriptions.
4.1 Text-Molecule Aligned Variational Autoencoder
Our text-molecule aligned variational autonecoder comprises of three parts: the molecular graph encoder $E_{g}$, text encoder $E_{t}$ and molecular graph decoder $D$. To bridge the representation gap between molecular graphs and their textual descriptions, the text and molecular graph encoders are trained on a large-scale dataset of molecule-text pairs. These encoders learn a well-aligned shared text-molecule representation, enriching the semantics of the molecular graph representation by aligning it the latent representation of textual descriptins of molecules through contrastive learning. These three parts are described in more detail below:
Molecular Graph Encoder aims to map each molecular graph $\mathcal{G}$ with the adjacency tensor $\mathbf{A}$ and node features matrix $\mathbf{X}$ into a lower-dimensional latent space $\mathcal{Z}$. This transformation maps the discrete space of graph structures and node attributes into a continuous latent space. Specifically, we use Graph Isomorphism Network (GIN) (Hu et al., 2019) as encoder network on the input adjacency tensor $\mathbf{A}$ and node features $\mathbf{X}$ of graph $\mathcal{G}$ to derive the mean and standard deviation of the variational marginals:
$\bm{\mu}_{g},\bm{\sigma}_{g}=\operatorname{GIN}(\mathbf{A},\mathbf{X}),$ (3)
where $\bm{\mu}_{g}$ and $\bm{\sigma}_{g}$ represent the mean and standard deviations, respectively. Then, the latent graph representation $\mathbf{z}$ is sampled from a Gaussian distribution as $\mathbf{z}\sim q\left(\mathbf{z}\mid\mathbf{A},\mathbf{X}\right)\simeq\mathcal{N}\left(\bm{\mu}_{g},\bm{\sigma}_{g}\right)$. This graph encoding process can also be expressed as $\mathbf{z}=E_{g}(\mathcal{G})$ signifying the outcome of the encoding operation within the input context of $\mathcal{G}$.
Text Encoder maps the textual descriptions of molecular graphs into a latent space, that provides input to a conditional generative model of molecular graphs. This encoded information becomes instrumental in generating molecules that match a given textual description. To inject potentially useful scientific knowledge from the literature into $E_{t}$ we initialize it with Sci-BERT (Beltagy et al., 2019), an encoder-only transformer pretrained on scientific publications.
Aligning the representation of molecular graphs with that of their textual descriptions is crucial for establishing a shared representation for effective text-guided molecular graph generation. Graph data consists of both node attributes and graph structure, which differs significantly from the sequential and contextual nature of text data. We aim to align the molecule representation from the molecule graph encoder $E_{g}$ with text representation from the text encoder $E_{t}$ by pretraining. Consider a molecular graph $\mathcal{G}$ and its textual description $\mathcal{T}$. To align these two encoders and ensure a cohesive latent space, we employ the following contrastive learning loss over a batch $\mathcal{B}\in\mathcal{D}$ of large-scale dataset containing molecule-text pairs to pretrain encoders:
$\mathcal{L}_{\text{con}}=-\frac{1}{|\mathcal{B}|}\sum_{(\mathcal{G}_{i},\mathcal{T}_{i})\in\mathcal{B}}\log\frac{\exp\left(\cos\left(\mathbf{z}_{i},\mathbf{c}_{i}\right)/\tau\right)}{\sum_{j=1}^{|\mathcal{B}|}\exp\left(\cos\left(\mathbf{z}_{i},\mathbf{c}_{j}\right)/\tau\right)},$ (4)
where $\mathbf{z}_{i}=E_{g}(\mathcal{G}_{i})$, $\mathbf{c}_{i}=E_{t}(\mathcal{T}_{i})$ and $\mathbf{c}_{j}=E_{t}(\mathcal{T}_{j})$ are representations of corresponding molecule graph and texts, respectively. $\cos(\cdot,\cdot)$ denotes cosine similarity and $\tau$ is the temperature hyperparameter. The encoders are pretrained on molecule-text (300K) pairs from PubChem collected by  Liu et al. (2023) and robust enough to capture a well-aligned molecule-text space, which overcomes the challenges posed by the distribution gap and produce more consistent and high-quality results in cross-modal molecule generation.
Molecular Graph Decoder learns to use latent representation $\mathbf{z}$ to decode the molecular graph. In this work, we employ the HierVAE (Jin et al., 2020) decoder, as it works empirically well in our final model. Note that our approach remains flexible and can be easily applied to other graph decoders (Li et al., 2018; Kong et al., 2022; Grover et al., 2019). Specifically, it generates a molecule $\mathcal{G}$ using a graph representation $\mathbf{z}$ from the encoder. Given the molecular graph $\mathcal{G}$, we minimize the negative Evidence Lower Bound (ELBO) (Jin et al., 2020) to train both encoder and decoder:
$\mathcal{L}_{\text{elbo}}=-\mathbb{E}_{q(\mathbf{z}\mid\mathbf{A},\mathbf{X})}[p(\mathcal{G}\mid\hat{\mathcal{G}})]+\alpha\mathrm{KL}[q(\mathbf{z}\mid\mathcal{G})\|p(\mathbf{z})].$ (5)
Here $\hat{\mathcal{G}}=D(\mathbf{z})$, in which $D$ denotes the decoder parameterized by HierVAE (Jin et al., 2020). $\mathrm{KL}[q(\cdot)||p(\cdot)]$ is the Kullback-Leibler (KL) divergence between $q(\textperiodcentered\cdot)$ and $p(\cdot)$. We follow Jin et al. (2020) and set the weight of KL loss $\alpha=0.1$. Here, $p(\mathbf{z})=\mathcal{N}\left(\mathbf{z}\mid\mathbf{0},\mathbf{I}\right)$ is the isotropic multivariate Gaussian prior, where $\mathbf{I}$ is the identity matrix. We optimize this first expectation term concurrently through reparameterization trick (Kingma & Welling, 2013).
4.2 Multi-modal Molecule Latent Diffusion
We then introduce latent diffusion for conditional molecular graph generation, a method that leverages the aligned molecule-text latent space of encoder-decoder discussed above, enabling us preserve the crucial molecular properties and aligning the textual description of $\mathcal{G}$. Specifically, we learn a better probabilistic mapping from the text to the aligned moelcule latent space, and thereby generate a higher-quality and more diverse molecules with more semantic consistency conforming to the conditional text inputs.
Specifically, the denoising network $\hat{\mathbf{z}}_{\theta}$ focuses on generating latent graph representation $\mathbf{z}$ conditioned on the text representation $\mathbf{c}$ from the text encoder $E_{t}$. The objective is:
$\small\mathcal{L}_{\text{diff}}=\mathbb{E}_{t,\mathbf{x}_{0},\epsilon}\left[\lambda_{t}\left\|\hat{\mathbf{z}}_{\theta}\left(\sqrt{\alpha_{t}}\mathbf{z}+\sqrt{1-\alpha_{t}}{\bm{\epsilon}},t,\mathbf{c}\right)-\mathbf{z}\right\|_{2}^{2}\right],$ (6)
where $\lambda_{t}$ are time-dependent weights and we use MLPs for the network $\hat{\mathbf{z}}_{\theta}$. Specifically, we concatenate c to graph latent representation $\mathbf{z}_{t}$ for timestep $t$ as the input of $\hat{\mathbf{z}}_{\theta}$. The denoising network is trained to denoise a noisy latent, $\mathbf{z}_{t}=\sqrt{\alpha_{t}}\mathbf{z}+\sqrt{1-\alpha_{t}}{\bm{\epsilon}}$ to the clean $\mathbf{z}$. $\mathbf{c}$ is the text representation from the fixed encoder $E_{t}$ pretrained by Equation (4).
In addition, we improve sample quality through classifier-free guidance (Ho & Salimans, 2022), jointly training an unconditional network, $\hat{\mathbf{z}}_{\theta}(\mathbf{z}_{t},t)$, and a conditional network, $\hat{\mathbf{z}}_{\theta}\left(\mathbf{z}_{t},t,\mathbf{c}\right)$. Conditioning information is randomly dropped with a probability of $p=0.1$ during training. When conditioning information is dropped, we replace the embedded source text with a learnable embedding.
4.3 Training and Inference
Training. We now describe the training procedure for 3M-Diffusion. While the learning objectives for graph encoder-decoder and multi-modal latent diffusion have already been specified by Equations (5) and (6), it is unclear whether the two components should be trained sequentially
Previous research on latent diffusion models for image generation indicates that a two-stage training strategy often results in superior performance (Rombach et al., 2022). Hence, we adopt a similar two-stage strategy. Specifically, our first stage is to train the encoder-decoder using Equation (5), and the second stage is to train multi-modal molecule latent diffusion with the fixed encoder-decoder from stage one.
Inference. During inference, we use a weight $w$ in computing the prediction as follows (Ho & Salimans, 2022):
$\tilde{\mathbf{z}}_{t}=w\hat{\mathbf{z}}_{\theta}\left(\mathbf{z}_{t},t,\mathbf{c}\right)+(1-w)\hat{\mathbf{z}}_{\theta}(\mathbf{z}_{t},t),$ (7)
where $\tilde{\mathbf{z}}_{t}$ represents the newly predicted results, incorporating both conditional and unconditional information. Setting $w=1.0$ corresponds to the conditional diffusion model, while setting $w>1.0$ enhances the impact of the conditioning information. A smaller value of $w$ contributes to more diverse samples. The new sampling equation is:
$\mathbf{z}_{t-1}$ $=\tfrac{\sqrt{\alpha_{t-1}}\left(1-\alpha_{t\mid t-1}\right)}{1-\alpha_{t}}\tilde{\mathbf{z}}_{t}$ (8)
$+\tfrac{\sqrt{\alpha_{t\mid t-1}}\left(1-\alpha_{t-1}\right)}{1-\alpha_{t}}\mathbf{z}_{t}+\sigma(t-1,t){\bm{\epsilon}},$
where the sampling chain is initialized from Gaussian prior $\mathbf{z}_{T}\sim p\left(z_{T}\right)=\mathcal{N}\left(\mathbf{z}_{T};\mathbf{0},\bm{I}\right)$. The detailed training and inference algorithms of 3M-diffusion are in Appendix B.
Table 1: Quantitative comparison of conditional generation on the ChEBI-20 and PubChem. 3M-Diffusion outperforms other SOTA methods in terms of novelty, diversity, and validity metrics by a large margin, while maintaining a good similarity metric.
PCDes MoMu
# Metrics Similarity (%) Novelty (%) Diversity (%) Validity (%) Similarity (%) Novelty (%) Diversity (%) Validity (%)
MolT5-small 64.84 24.91 9.67 73.96 16.64 97.49 29.95 60.19
MolT5-base 71.71 25.85 10.50 81.92 19.76 97.78 29.98 68.84
MolT5-large 88.37 20.15 9.49 96.48 25.07 97.47 30.33 90.40
ChemT5-small 86.27 23.28 13.17 93.73 23.25 96.97 30.04 88.45
ChemT5-base 85.01 25.55 14.08 92.93 23.40 97.65 30.07 87.61
3M-Diffusion 81.57 63.66 32.39 100.0 24.62 98.16 37.65 100.0
Table 2: Quantitative comparison of conditional generation on the ChEBI-20 and PubChem. 3M-Diffusion outperforms other SOTA methods in terms of novelty, diversity, and validity metrics by a large margin, while maintaining a good similarity metric.
ChEBI-20 PubChem
# Methods Similarity (%) Novelty (%) Diversity (%) Validity (%) Similarity (%) Novelty (%) Diversity (%) Validity (%)
MolT5-small 73.32 31.43 17.22 78.27 68.36 20.63 9.32 78.86
MolT5-base 80.75 32.83 17.66 84.63 73.85 21.86 9.89 79.88
MolT5-large 96.88 12.92 11.20 98.06 91.57 20.85 9.84 95.18
ChemT5-small 96.22 13.94 13.50 96.74 89.32 20.89 13.10 93.47
ChemT5-base 95.48 15.12 13.91 97.15 89.42 22.40 13.98 92.43
3M-Diffusion 87.09 55.36 34.03 100 87.05 64.41 33.44 100
5 Experiment
5.1 Experimental Settings
Table 3: Statistics of all datasets.
Dataset #Training #Validation #Test
ChEBI-20 15,409 1,971 1,965
PubChem 6,912 571 1,162
PCDes 7,474 1,051 2,136
MoMu 7,474 1,051 4,554
Datasets. In our experiments, we use PubChem (Liu et al., 2023), ChEBI-20 (Edwards et al., 2021), PCDes (Zeng et al., 2022) and Momu (Su et al., 2022) datasets. Following the evaluation methodology adopted with datasets (ZINC250K (Irwin et al., 2012) and QM9 (Blum & Reymond, 2009; Rupp et al., 2012)) used in previous studies of automated generation of molecular structures, we limit our training, validation, and test data to molecules with fewer than 30 atoms. Statistics of datasets are shown in Table 3 and detailed descriptions are given in the Appendix C.1.
Evaluation. We evaluate the model’s performance on two tasks: text-guided and unconditional molecule generation. Following previous works on text-guided molecule generation, we compare our 3M-Diffusion the SOTA baselines (Edwards et al., 2022; Christofidellis et al., 2023). Text-guided molecule generation measures the model’s capacity to generate chemically valid, structurally diverse, novel molecules that match the textual description provided. Based on previous work (Chen et al., 2021; Fu et al., 2021; Wang et al., 2022), we adopt the following metrics for this task: Similarity, Novelty, Diversity, Validity. We call a predicted molecular graph, specifically one that is chemically valid ($g^{\prime}$), qualified with the ground truth graph $g$ if $f(g,g^{\prime})>0.5$, where $f$ measures the MACCS (Durant et al., 2002) cosine similarity between the two. If the $f(g,g^{\prime})$ similarity between $g^{\prime}$ and $g$ is smaller than a threshold (0.8 in this paper), we call this predicted molecule $g$ novel. Specifically, (i) Similarity: the percentage of qualified molecules out of 20K molecules; (ii) Novelty: the percentage of novel molecules out of all qualified molecules; (iii) Diversity: the average pairwise distance $1-f(\cdot,\cdot)$ between all qualified molecules. (iv) Validity: the percentage of generated molecules that are chemically valid out of all molecules.
We report the results for unconditional molecule generation tasks using the GuacaMol benchmarks (Brown et al., 2019) to assess the model’s capability to learn molecular data distribution and generate realistic, diverse molecules. Specifically, Uniqueness gauges the ratio of unique molecules among the generated ones. Novelty measures the models’ capacity to generate molecules not present in the training set. KL Divergence assesses the proximity between the distributions of various physicochemical properties in the training set and the generated molecules. Fréchet ChemNet Distance (FCD) computes the proximity of two sets of molecules based on their hidden representations in ChemNet (Preuer et al., 2018). Each metric is normalized to a range of 0 to 1, with a higher value indicating better performance.
Table 4: Quantitative comparison of unconditional generation. Results of Uniq, KL Div and FCD on ChEBI-20 and PubChem, which refer to Uniqueness, KL Divergence and Fréchet ChemNet Distance, respectively. A higher number indicates a better generation quality.
ChEBI-20 PubChem
# Methods Uniq (%) Novelty (%) KL Div (%) FCD (%) Validity (%) Uniq (%) Novelty (%) KL Div (%) FCD (%) Validity (%)
BwR 22.09 21.97 50.59 0.26 22.66 82.35 82.34 45.53 0.11 87.73
HierVAE 82.17 72.83 93.39 64.32 100.0 75.33 72.44 89.05 50.04 100.0
PS-VAE 76.09 74.55 83.16 32.44 100.0 66.97 66.52 83.41 14.41 100.0
3M-Diffusion 83.04 70.80 96.29 77.83 100.0 85.42 81.20 92.67 58.27 100.0
Baselines. For text guided molecule generation, we compare 3M-Diffusion with the following baselines: MolT5 (Edwards et al., 2022) and ChemT5 (Christofidellis et al., 2023). We consider three representative baselines for unconditional molecule generation based on Variational Autoencoder, including BwR (Diamant et al., 2023), HierVAE (Jin et al., 2020) and PS-VAE (Kong et al., 2022). The descriptions and details of the baselines are given in Appendix C.1.
Setup. We randomly initialize the model parameters and use the Adam optimizer to train the model. HierVAE is used as our variational autoencoder for all of the following experiments. For baselines of unconditional generation that did not report results in these datasets, we reproduce the results using the official code of the authors. To ensure a fair comparison across all methods for unconditional generation, we select the best hyperparameter configuration solely based on the loss on the training set. We evaluate the performance of baselines for text-guided molecule generation using the released pre-trained parameters.
Figure 3: Molecules generated unconditionally by 3M-diffusion from the prior distribution $\mathcal{N}(\mathbf{0},\bm{I})$.
5.2 Text-guided Molecule Generation
Table 1 and 2 report Similarity, Novelty, Diversity, and Validity for text-guided molecule generation on four datasets. All of these metrics are computed as averages, with each metric being calculated based on 5 samples generated from each textual description. Our model demonstrates competitive results in Similarity and notably higher values in the Diversity and Novelty. Specifically, 3M-Diffusion achieves 146.27% (Novelty of PCDes), 130.04% (Diversity of PCDes) relative improvement over the second-best method with maintaining semantics in the textual prompt. This suggests that transformer-based models, pre-trained on a large dataset, excel in predicting molecules based on textual descriptions. However, our model can generate a diverse and novel collection of molecular graphs that match the textual descriptions provided as input. Also, as Figure  4 shows, our model can generate molecules faster than transformer-based methods.
5.3 Unconditional Molecule Generation
Table 4 shows the results of unconditional molecule generation on PubChem and ChEBI-20. 3M-Diffusion consistently outperforms or achieves competitive results across all four metrics on both datasets, indicating its ability to generate realistic molecules while avoiding overfitting the training data. Furthermore, 3M-Diffusion consistently outperforms HierVAE when both methods use the same decoder. This underscores the clear advantage of 3M-Diffusion in generating a diverse and novel assortment of molecular graphs. As shown in Figure 3, our sampled molecules present rich variety and structural complexity. We have included more sampled molecules from the prior distribution in Appendix D.4.
Figure 4: The average inference time comparison of molecule conditional generation on ChEBI-20 dataset.
5.4 Ablation Study
We conduct ablation studies to assess the impact of different model designs, including joint training of the diffusion model and Variational Autoencoder (denoted as ”Without two stage training”) and training the model without aligning the representation space of graphs and text using $\mathcal{L}_{\text{con}}$ in Equation (4) (denoted as ”Without representation alignment”). Additionally, we perform experiments without using either of these configurations, labeled as ”Without two stage training & Without representation alignment” in Table 5. Initially, training the aligned encoder-decoder alongside Multi-modal Molecule Latent Diffusion jointly yielded comparable or superior performance in terms of diversity and novelty. However, there was a noticeable deficiency in its ability to learn align its representation of molecular graphs with that of their textual descriptions. Training 3M-Diffusion without the aligned encoder-decoder results in substantial drop in performance relative to 3M-Diffusion. This result underscores the importance of aligning the representation of textual descriptions of molecules with that of molecular graphs. Without using both design, Similarity suffers a substantial decline to 10.24, rendering it unsuitable for real-world applications. The full 3M-Diffusion model (last row) achieves the best performance with respect to all of the metrics. This confirms the suitability of the two-step training approach for generating diverse, novel set of molecules based on their textual descriptions.
Table 5: Results of ablation study with different model designs of 3M-diffusion for the Pubchem dataset.
# Methods Similarity (%) Novelty (%) Diversity (%) Validity (%)
(i) Without two-stage training 77.74 67.03 31.05 100.0
(ii) Without representation alignment 86.21 62.81 32.78 100.0
(iii) Without two-stage training & representation alignment 10.24 77.89 27.86 100.0
3M-Diffusion 87.05 64.41 33.44 100.0
5.5 Qualitative Analysis and Case Study
Figure 5: Molecules generated conditionally on input text by 3M-Diffusion on the ChEBI-20 dataset.
Property Conditional Molecule Generation. In traditional molecule generation models, the generation process is often conditioned on specific molecular properties such as logP (Kusner et al., 2017) and QED. logP assesses a molecule’s solubility, while QED evaluates its drug likeness (Bickerton et al., 2012). Previous methods have explored ways to optimize their models further in order to generate molecules with desired properties. However, in the case of our model, a different approach is taken. Our model utilizes designed prompts to realize the conditional molecule generation process, as illustrated in Figure 2 for assessing molecule solubility. In particular, we generated 10 samples for each prompt and subsequently select the top 5 samples based on the desired property values. For MolT5-large, we adjust the temperature of the transformer in an attempt to encourage diversity while ensuring that invalid samples are controlled during the sampling process. We applied the same methodology used in our previous experiments to maintain consistency across the comparative analysis, as shown in Figure 2. Compared with the MolT5-large model, we find that 3M-Diffusion excels in generating a diverse range of molecules while achieving improved desired properties, as evidenced by higher logP values Figure 2. This underscores our model’s capacity to produce molecules that are not only novel and diverse but also match the textual description provided. This unique approach distinguishes 3M-Diffusion from prior methods for molecular generation, presenting a fresh and promising perspective on molecular design. For additional experimental details and insights regarding prompts related to solubility and drug likeness, please refer to the Appendix D.1.
Qualitative Comparison. In Figure 2, we introduce more specific prompts related to molecular categories, such as ”The molecule is a member of the class of flavones.” These prompts bias the model, in favor of generating useful drug candidates. Specifically, we choose a reference molecule from the training dataset that matches the given textual prompt. Subsequently, we generate ten additional molecules and select the top five molecules that exhibit the highest structural similarity (calculated by Rascal algorithm (Raymond et al., 2002)) to the reference molecule. These selected molecules are shown in Figure 2.Comparisong of our model with MolT5-large shows that our model generates more diverse molecules that exhibit a higher degree of structural similarity to the reference molecule. Simultaneously, it also produces novel molecular graphs. We also include more prompt examples in Appendix D.2.
Case Study. In Figure 5, we choose a set of examples from the ChEBI-20 dataset to visually illustrate the structural similarity between the molecules generated by 3M-Diffusion and the ground truth molecules corresponding to the given text descriptions. Specifically, we show the largest substructure that appears in generated and ground truth molecules by RDKit. 3M-Diffusion demonstrates the capability to generate molecules that share a common substructure, as illustrated in Figure 5. Simultaneously, it has the flexibility to explore novel components or structures, thereby achieving a balance between generating similar structures and introducing novel molecules. This observation further confirms the ability of our model in generating a diverse collection of novel molecular structures that match the textual description provided. We show more results in Appendix D.3.
6 Conclusion
Generating molecules with desired properties is a critical task with broad applications in drug discovery and materials design. These applications call for methods that generate diverse, and ideally novel, molecules with the desired properties. We have introduced 3M-Diffusion, a novel multi-modal molecular graph generation method, to address this challenge. 3M-Diffusion first encodes molecular graphs into a graph latent space aligned with text descriptions. It then reconstructs the molecular graph and atomic attributes based on the given text descriptions using the molecule decoder. 3M-Diffusion learns a probabilistic mapping from the text space to the latent molecular graph space using the latent diffusion model. The results of our extensive experiments demonstrate that 3M-Diffusion can generate high-quality, novel and diverse molecular graphs that semantically match the textual description provided. Ablation studies underscore the importance of aligning the latent representation of texts with that of molecular graphs to the performance of 3M-Diffusion. Work in progress is aimed at text-guided molecular graph generation tasks, e.g., text-guided generation of biological macromolecules, new drugs and new materials, and experimental validation of the resulting graphs through their synthesis and characterization in the lab.
Appendix A Diffusion Models
We present a formal description of diffusion (Ho et al., 2020; Song et al., 2020). Diffusion models are latent variable models that represent the data $\mathbf{z}_{0}$ as Markov chains $\mathbf{z}_{T}\cdots\mathbf{z}_{0}$, where intermediate variables share the same dimension. DMs involve a forward diffusion process $q\left(\mathbf{z}_{1:T}\mid\mathbf{z}_{0}\right)=\prod_{t=1}^{T}q\left(\mathbf{z}_{t}\mid\mathbf{z}_{t-1}\right)$ and a reverse process $p_{\theta}\left(\mathbf{z}_{0:T}\right)=p\left(\mathbf{z}_{T}\right)\prod_{t=1}^{T}p_{\theta}\left(\mathbf{z}_{t-1}\mid\mathbf{z}_{t}\right)$. The forward process starting from $\mathbf{z}_{0}$ can be written as:
$q\left(\mathbf{z}_{t}\mid\bm{x}_{0}\right):=\int q\left(\mathbf{z}_{1:t}\mid\bm{x}_{0}\right)\mathrm{d}\mathbf{z}_{1:(t-1)}=\mathcal{N}\left(\mathbf{z}_{t};\sqrt{\alpha_{t}}\bm{x}_{0},\left(1-\alpha_{t}\right)\mathbf{I}\right),$ (9)
where the hyperparameter $\alpha_{1:T}$ determines the magnitude of noise added at each timestep $t$. The values of $\alpha_{1:T}$ are selected to ensure that samples $\mathbf{z}_{T}$ converge to standard Gaussians, i.e., $q\left(\mathbf{z}_{T}\right)\approx\mathcal{N}(0,\bm{I})$. This forward process $q$ is usually predefined without trainable parameters.
The generation process of DMs is defined as learning a parameterized reverse denoising process, which incrementally denoises the noisy variables $\mathbf{z}_{T:1}$ to approximate clean data $x_{0}$ in the target data distribution:
$p_{\theta}\left(\mathbf{z}_{t-1}\mid\mathbf{z}_{t}\right)=\mathcal{N}\left(\mathbf{z}_{t-1};\bm{\mu}_{\theta}\left(\mathbf{z}_{t},t-1,t\right),\sigma^{2}(t-1,t)\bm{I}\right),$ (10)
where the initial distribution $p(\mathbf{z}_{T})$ is defined as $\mathcal{N}(0,\bm{I})$. The means $\mu_{\theta}$ typically are neural networks such as U-Nets for images or Transformers for text, and the variances $\sigma^{2}(t-1,t)$ typically are also predefined. As latent variable models, the forward process $q(\mathbf{z}_{1:T}|\mathbf{z}_{0})$ can be seen as a fixed posterior. The reverse process $p_{\theta}(\mathbf{z}_{0:T})$ is trained to maximize the variational lower bound of the likelihood of the data:
$\mathcal{L}_{\text{vlb}}=\mathbb{E}_{q\left(\mathbf{z}_{1:T}\mid\mathbf{z}_{0}\right)}\left[\log\frac{q\left(\mathbf{z}_{T}\mid\mathbf{z}_{0}\right)}{p_{\theta}\left(\mathbf{z}_{T}\right)}+\right.\left.\sum_{t=2}^{T}\log\frac{q\left(\mathbf{z}_{t-1}\mid\mathbf{z}_{0},\mathbf{z}_{t}\right)}{p_{\theta}\left(\mathbf{z}_{t-1}\mid\mathbf{z}_{t}\right)}-\log p_{\theta}\left(\mathbf{z}_{0}\mid\mathbf{z}_{1}\right)\right]$ (11)
To ensure training stability, a simple surrogate objective up to irrelevant constant terms is introduced (Song et al., 2020; Ho et al., 2020; Nichol & Dhariwal, 2021):
$\mathcal{L}_{\text{diff}}=\mathbb{E}_{t,\mathbf{z}_{0},\epsilon}\left[\lambda_{t}\left\|\hat{\mathbf{z}}_{\theta}\left(\sqrt{\alpha_{t}}\mathbf{z}_{0}+\sqrt{1-\alpha_{t}}{\bm{\epsilon}},t\right)-\mathbf{z}_{0}\right\|_{2}^{2}\right],$ (12)
where $\mathbf{z}_{t}$ is used to denote $\sqrt{\alpha_{t}}\mathbf{z}_{0}+\sqrt{1-\alpha_{t}}{\bm{\epsilon}}$ for simplicity. Intuitively, we train a neural network $\hat{\mathbf{z}}_{\theta}(\mathbf{z}_{t},t)$ to approximate the original data given some noisy latent and the timestep through Equation (12), $\hat{\mathbf{z}}_{\theta}\left(\mathbf{z}_{t},t\right)\approx\mathbf{z}$. With a trained denoising network, we define the mean and variance function in Equation (10) following the previous work (Ho et al., 2020) as:
$\bm{\mu}_{\theta}\left(\mathbf{z}_{t},t-1,t\right)=\frac{\sqrt{\alpha_{t-1}}\left(1-\alpha_{t\mid t-1}\right)}{1-\alpha_{t}}\hat{\mathbf{z}}_{\theta}(\mathbf{z}_{t},t)+\frac{\sqrt{\alpha_{t\mid t-1}}\left(1-\alpha_{t-1}\right)}{1-\alpha_{t}}\mathbf{z}_{t},\sigma^{2}(t-1,t)=\frac{\left(1-\alpha_{t-1}\right)\left(1-\alpha_{t\mid t-1}\right)}{1-\alpha_{t}},$ (13)
where $\lambda_{t}$ is the time-dependent weight and $\alpha_{t\mid t-1}=\alpha_{t}/\alpha_{t-1}$. For generation, we employ the standard DDPM sampler, also referred to as the ancestral sampler (Ho et al., 2020). The process involves sampling initial noise $\mathbf{z}_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ and iteratively applying the update rule:
$\mathbf{z}_{t-1}$ $=\frac{\sqrt{\alpha_{t-1}}\left(1-\alpha_{t\mid t-1}\right)}{1-\alpha_{t}}\hat{\mathbf{z}}_{\theta}(\mathbf{z}_{t},t)+\frac{\sqrt{\alpha_{t\mid t-1}}\left(1-\alpha_{t-1}\right)}{1-\alpha_{t}}\mathbf{z}_{t}+\sigma(t-1,t){\bm{\epsilon}},$ (14)
where ${\bm{\epsilon}}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. We use $T=50$ for the sampling timesteps.
Appendix B Training and Sampling Algorithm
The complete training algorithm is depicted in Algorithm 1. Firstly, the process begins by training the graph and text encoders using contrastive learning. Following that, the Variational Autoencoder is trained to facilitate the reconstruction of molecule graphs. Lastly, the Multi-modal Molecule Latent Diffusion Model is employed to train the denoising network for the sampling process. This structured approach forms the core of our training procedure. Moreover, we present the sampling method of 3M-Diffusion in Algorithm 2.
Algorithm 1 Training Algorithm of 3M-Diffusion
1:  Input: Pairs of Molecule Graph data $\mathcal{G}=(\mathcal{V},\mathcal{E})$ and Textual Description $\mathcal{T}$
2:  Initial: Graph Encoder network $E_{g}$, Text Encoder network $E_{t}$ and Decoder network $D$, denoising network $\hat{\mathbf{z}}_{\theta}$
3:  Pretraining Stage: Aligned Text and Graph Encoder Training
4:  while $E_{g}$ and $E_{t}$ have not converged do
5:     $\mathcal{L}_{\text{con}}=-\frac{1}{|\mathcal{B}|}\sum_{(\mathcal{G}_{i},\mathcal{T}_{i})\in\mathcal{B}}\log\frac{\exp\left(\cos\left(\mathbf{z}_{i},\mathbf{c}_{i}\right)/\tau\right)}{\sum_{j=1}^{|\mathcal{B}|}\exp\left(\cos\left(\mathbf{z}_{i},\mathbf{c}_{j}\right)/\tau\right)}$
6:     $E_{g},E_{t}\leftarrow\operatorname{optimizer}(\mathcal{L}_{\text{con}}$)
7:  end while
8:  First Stage: Autoencoder Training
9:  Initialize $E_{g}$ from Pretraining Stage
10:  while $E_{g}$ and $D$ have not converged do
11:     $\mathbf{z}\leftarrow E_{g}(\mathcal{G})$
12:     $\mathcal{G}\leftarrow D(\mathbf{z})$blacktriangleright Decoding
13:     $\mathcal{L}_{\text{elbo}}=-\mathbb{E}_{q(\mathbf{z}\mid\mathbf{A},\mathbf{X})}[p(\mathcal{G}\mid\hat{\mathcal{G}})]+\alpha\mathrm{KL}[q(\mathbf{z}\mid\mathcal{G})\|p(\mathbf{z})]$
14:     $E_{g},D\leftarrow\operatorname{optimizer}(\mathcal{L}_{\text{elbo}})$
15:  end while
16:  Second Stage: Multi-modal Molecule Latent Diffusion Training
17:  Fix Graph Encoder $E_{g}$ from Stage 1 and Text Encoder $E_{t}$ from Pretraining Stage
18:  while $\mathbf{z}_{\theta}$ have not converged do
19:     $\mathbf{z}\leftarrow E_{g}(\mathcal{G})$
20:     $t\sim{\mathbf{U}}(0,T)$, ${\bm{\epsilon}}\sim{\mathcal{N}}(\bm{0},{\bm{I}})$
21:     $\mathcal{L}_{\text{diff}}=\mathbb{E}_{t,\mathbf{x}_{0},\epsilon}\left[\lambda_{t}\left\|\hat{\mathbf{z}}_{\theta}\left(\sqrt{\alpha_{t}}\mathbf{z}_{0}+\sqrt{1-\alpha_{t}}{\bm{\epsilon}},t\right)-\mathbf{z}_{0}\right\|_{2}^{2}\right]$
22:     $\theta\leftarrow\operatorname{optimizer}(\mathcal{L}_{\text{diff}})$
23:  end while
24:  return $E_{g}$, $D$
Algorithm 2 Sampling Algorithm of 3M-Diffusion
1:  Input: decoder network $D$, denoising network $\hat{\mathbf{z}}_{\theta}$
2:  $\mathbf{z}_{T}\sim{\mathcal{N}}(\mathbf{0},{\bm{I}})$
3:  for $t$ in $T,\,T-1,\cdots,1$ do
4:     ${\bm{\epsilon}}\sim\mathcal{N}(\mathbf{0},{\bm{I}})$blacktriangleright Latent Denoising Loop
5:     $\tilde{\mathbf{z}}_{t}=w\hat{\mathbf{z}}_{\theta}\left(\mathbf{z}_{t},t,\mathbf{c}\right)+(1-w)\hat{\mathbf{z}}_{\theta}(\mathbf{z}_{t},t),$
6:     $\mathbf{z}_{t-1}=\tfrac{\sqrt{\alpha_{t-1}}\left(1-\alpha_{t\mid t-1}\right)}{1-\alpha_{t}}\tilde{\mathbf{z}}_{t}+\tfrac{\sqrt{\alpha_{t\mid t-1}}\left(1-\alpha_{t-1}\right)}{1-\alpha_{t}}\mathbf{z}_{t}+\sigma(t-1,t){\bm{\epsilon}}$
7:  end for
8:  $\mathcal{G}\sim p(\mathcal{G}|\mathbf{z}_{0})$blacktriangleright Decoding
9:  return ${\mathbf{x}},{\mathbf{h}}$
Appendix C Experimental Details
C.1 Datasets Details and Statistics
The statistics of datasets, including their the number of samples for training, validation and test sets are given in Table 3.
ChEBI-20 (Edwards et al., 2021): This dataset is generated by leveraging PubChem (Kim et al., 2016) and Chemical Entities of Biological Interest (ChEBI) (Hastings et al., 2016). It compiles ChEBI annotations of compounds extracted from PubChem, comprising molecule-description pairs. To ensure less noise and more informative molecule descriptions, it includes samples with descriptions exceeding 20 words.
PubChem (Liu et al., 2023): This dataset comprises 324k molecule-text pairs gathered from the PubChem website. As the dataset contains numerous uninformative texts like ”The molecule is a peptide,” it curates a high-quality subset of 15k pairs with text longer than 19 words for downstream tasks. This refined subset is then randomly split into training, validation, and test sets. The remaining dataset, which is more noisy, is utilized for pretraining. To prevent data leakage, we exclude samples that appear in the testing set of other datasets from the pretraining dataset.
PCDes (Zeng et al., 2022): This dataset compiles molecule-text pairs sourced from PubChem (Kim et al., 2016), encompassing names, SMILES notations, and accompanying paragraphs providing property descriptions for each molecule.
MoMu (Su et al., 2022): This dataset comprises paired molecule graph-text data, with the textual information for each molecule extracted from the SCI paper dataset (Lo et al., 2019).
For all datasets, inspired by previous datasets (ZINC250K (Irwin et al., 2012) and QM9 (Blum & Reymond, 2009; Rupp et al., 2012)) on molecule generation, we preserve molecules with fewer than 30 atoms, creating new training, validation, and test sets from their original counterparts.
C.2 Baselines
We firstly introduce the baselines used for text-guided molecule generation in this section.
MolT5 (Edwards et al., 2022): MolT5 addresses the difficult problem of cross-domain generation by linking natural language and chemistry, tackling tasks such as text-conditional molecule generation and molecule captioning. It presents several T5-based LMs (Raffel et al., 2020) by training on the SMILES-to-text and text-to-SMILES translations.
ChemT5 (Christofidellis et al., 2023): ChemT5 introduces a versatile multi-task model for seamless translation between textual and chemical domains. This innovative approach leverages transfer learning in the chemical domain, with a particular emphasis on addressing cross-domain tasks that involve both chemistry and natural language.
Then, we will also include three baselines for unconditional generation to verify the effectiveness of our model.
BwR (Diamant et al., 2023): BwR aims to decrease the time/space complexity and output space of graph generative models, demonstrating notable performance improvements in generating high-quality outputs on biological and chemical datasets.
HierVAE (Jin et al., 2020): HierVAE employs a hierarchical encoder-decoder architecture to generate molecular graphs, leveraging structural motifs as fundamental building blocks.
PS-VAE (Kong et al., 2022): PS-VAE automatically identifies regularities within molecules, extracting them as principal subgraphs. Utilizing these extracted principal subgraphs, it proceeds to generate molecules in two distinct phases.
C.3 Setup and Hyper-parameter settings
We use official implementation publicly released by the authors of the baselines or implementation from Pytorch. Note that we use the decoder of HierVAE for our model. We run experiments on a machine with a NVIDIA A100 GPU with 80GB of GPU memory. In all experiments, we use the Adam optimizer (Kingma & Ba, 2014). In particular, for our 3M-diffusion. we set the dimension of latent representation as 24, the number of layers for the denosing networks (MLP) as 4, $\alpha$ as 0.1, $T$ as 100 for training and 50 for sampling process and the probability $p=0.8$ for text-guided molecule generation. For unconditional generation, we set $T$ as 10 or 5 for the sampling process.
Appendix D More Visualization Results
D.1 Property Conditional Molecule Generation
In this section, we provide more prompt examples and more experiment details for Property Conditional Molecule Generation. We generate 10 samples for each prompt and subsequently select the top 5 samples based on the desired property values, as shown in Figure 2,  6,  7 and  8. For instance, when using the prompt ”This molecule is not like a drug,” we select the top 5 molecules with the smallest QED values, as indicated. Conversely, for the prompt ”This molecule is like a drug,” we opt for the top 5 molecules with the largest QED values. This same approach is applied to solubility, where we select molecules based on logP values. Specifically, we adjust temperature setting within the MolT5 transformer model to promote the generation of diverse molecules. We carefully selected an appropriate temperature threshold that struck a balance between diversity and maintaining the validity of the generated molecules. Moreover, when dealing with MolT5, it is worth noting that it may occasionally generate multiple unconnected molecules within a single SMILES representation. In such cases, we select the molecule among these unconnected ones that best aligns with the top desired attribute. From all figures, it becomes evident that our model excels in generating molecules with improved desired properties when compared to MolT5-large.
Figure 6: Qualitative comparisons to the MolT5-large in terms of generated molecules on CheBI-20. Compared with the SOTA method MolT5-large, our generated results are more diverse and novel with maintained semantics in textual prompt.
Figure 7: Qualitative comparisons to the MolT5-large in terms of generated molecules on CheBI-20. Compared with the SOTA method MolT5-large, our generated results are more diverse and novel with maintained semantics in textual prompt.
Figure 8: Qualitative comparisons to the MolT5-large in terms of generated molecules on CheBI-20. Compared with the SOTA method MolT5-large, our generated results are more diverse and novel with maintained semantics in textual prompt.
D.2 Qualitative Comparison
Figure 9: Qualitative comparisons to the MolT5-large in terms of generated molecules on CheBI-20. Compared with the SOTA method MolT5-large, our generated results are more diverse and novel with maintained semantics in textual prompt.
In this section, we introduce an additional prompt example: ”The molecule is an anthocyanidin cation.” This prompt is used to generate molecules belonging to specific categories. We have the same observation that our model excels in generating molecules that not only closely resemble the reference molecule but also encompass some diverse and novel structures.
D.3 Case Study
In this section, we choose 20 more examples from the ChEBI-20 dataset to visually illustrate the structural similarity between the molecules generated by 3M-Diffusion and the ground truth molecules corresponding to the given text descriptions. All results are shown in Figure 5. We also observe that our model is proficient at generating molecules that share significant common structural elements, while simultaneously obtaining diverse and innovative molecular structures.
Figure 10: More molecules generated conditionally on input text by 3M-Diffusion on the ChEBI-20 dataset.
D.4 Unconditional Molecule Generation
We further present 50 molecules sampled from the prior distribution in Figure 11. As shown in Figure 11, our sampled molecules present rich variety and structural complexity. This demonstrates our model’s ability to generate diverse, novel, and realistic molecules.
Figure 11: 50 molecules sampled from the prior distribution $\mathcal{N}(\mathbf{0},\bm{I})$.

### v2

@@
-3M-Diffusion: Latent Multi-Modal Diffusion for Text-Guided Generation of Molecular Graphs
+3M-Diffusion: Latent Multi-Modal Diffusion for Language-Guided Molecular Structure Generation
 Abstract
-Generating molecules with desired properties is a critical task with broad applications in drug discovery and materials design.
-Inspired by recent advances in large language models, there is a growing interest in using natural language descriptions of molecules to generate molecules with the desired properties.
-Most existing methods focus on generating molecules that precisely match the text description.
-However, practical applications call for methods that generate diverse, and ideally novel, molecules with the desired properties.
-We propose 3M-Diffusion, a novel multi-modal molecular graph generation method, to address this challenge.
-3M-Diffusion first encodes molecular graphs into a graph latent space aligned with text descriptions.
+Generating molecular structures with desired properties is a critical task with broad applications in drug discovery and materials design.
+We propose 3M-Diffusion, a novel multi-modal molecular graph generation method, to generate diverse, ideally novel molecular structures with desired properties.
+3M-Diffusion encodes molecular graphs into a graph latent space which it then aligns with the text space learned by encoder-based LLMs from textual descriptions.
 It then reconstructs the molecular structure and atomic attributes based on the given text descriptions using the molecule decoder.
@@
 The results of our extensive experiments on several datasets demonstrate that 3M-Diffusion can generate high-quality, novel and diverse molecular graphs that semantically match the textual description provided.
-Code is released at https://github.com/huaishengzhu/3MDiffusion.
-Machine Learning, ICML
+The code is available on github.
+**footnotetext: Equal contribution
 1 Introduction
-Generating molecules with the desired properties is a critical task with broad applications in drug discovery and materials design (Hajduk & Greer, 2007; Mandal et al., 2009; Pyzer-Knapp et al., 2015).
-There is a growing interest in generative models to automate this task (You et al., 2018; Jin et al., 2018, 2020; Bjerrum & Threlfall, 2017).
+Generating molecular structures with the desired properties is a critical task with broad applications in drug discovery and materials design (Hajduk & Greer, 2007; Mandal et al., 2009; Pyzer-Knapp et al., 2015).
+There is a growing interest in generative models to automate this task (You et al., 2018; Jin et al., 2018; 2020; Bjerrum & Threlfall, 2017).
 Kusner et al.
@@
 (2021) formulated molecule retrieval from textual description as a cross-lingual retrieval task.
-Recent advances in language models have led to innovative approaches to generate molecules from human-friendly textual descriptions of their properties, substructures, and biochemical activity (Edwards et al., 2022; Zeng et al., 2022; Zhao et al., 2023b; Fang et al., 2023a).
-These approaches learn mappings from textual descriptions to chemical structures using single transformer-based cross-lingual language model for molecule generation.
-Specifically, with the given text description of the molecule, these approaches produce SMILES strings describing the molecular graph.
-Figure 1: The overview of 3M-diffusion, with a molecule graph encoder/decoder and a latent diffusion model conditioned on a prior (an aligned text encoder).
+Recent advances in language models have led to innovative approaches to generate molecular structures that match human-friendly textual descriptions of their properties, substructures, and biochemical activity (Edwards et al., 2022; Zeng et al., 2022; Zhao et al., 2023b; Fang et al., 2023a).
+These approaches learn mappings from textual descriptions to molecular structures using single transformer-based cross-lingual language model for molecule generation.
+Specifically, with the given text description of a molecule, these approaches produce SMILES strings that encode the molecular structure.
+Despite promising initial results, such approaches suffer from some important limitations: (i) SMILES encoding is degenerate in that the a given molecular structure can have multiple SMILE encodings.
+For example, the structure of ethanol can be specified using CCO, OCC or C(O)C.
+Although in principle a unique canonical SMILES representation can be produced for a given molecular structure, the procedure used to do so relies on several arbitrary choices.
+Not surprisingly, molecules with identical structure may map to different SMILES strings.
+This makes SMILES notation less than ideal for generating molecules from their textual descriptions (Jin et al., 2018).
+Important chemical properties corresponding to the substructures of molecules are more straightforward to express using molecular graphs rather than linear SMILES representations (Jin et al., 2018; Du et al., 2022); (ii) Using language models to generate diverse and high-quality molecular structures matching a description is challenging, because the process used to decode the learned representation into the corresponding molecular structure, e.g., beam search or greedy search, often produces structures that are often too similar to each other (See Figure 2).
+Hence, there is an urgent need for better approaches for generating diverse, novel molecular structures from a given textual description  (Brown et al., 2019).
+Against this background, this paper aims to answer the following research question: How can we generate a diverse and novel collection of molecular graphs that match a given textual description?
+Figure 1: The overview of 3M-diffusion, with a molecular graph encoder/decoder and a latent diffusion model conditioned on a prior (an aligned LLM encoder).
 The details of alignment in the text/graph encoders and diffusion model are given in Section 4.
-Figure 2: Qualitative comparisons to the MolT5-large in terms of generated molecules on ChEBI-20.
-Compared with the SOTA method MolT5-large, our generated results are more diverse and novel with maintained semantics in textual prompt.
-Despite promising initial results, such approaches suffer from some important limitations: (i) SMILES encoding is degenerate in that the a given molecular graph can have multiple SMILE encodings.
-For example, the structure of ethanol can be specified using CCO, OCC or C(O)C.
-Although a unique canonical SMILES representation can be produced for a given graph, the procedure used to do so relies on several arbitrary choices.
-Not surprisingly, molecules with identical chemical graphs may map to significantly distinct SMILES strings, making SMILES notation less than ideal for generating molecules from their textual descriptions (Jin et al., 2018).
-Important chemical properties corresponding to the substructures of molecules are more straightforward to express using molecular graphs rather than linear SMILES representations (Jin et al., 2018; Du et al., 2022); (ii) Using language models to generate diverse high-quality molecules matching a description is challenging, because the decoding process used such as beam search or greedy search often produces candidates that are often too similar to each other as shown in Figure 2.
-Hence, there is an urgent need for better approaches to generating diverse, novel molecular graphs from a given textual description  (Brown et al., 2019).
-Against this background, this paper aims to answer the following research question: How can we generate a diverse and novel collection of molecular graphs that match a given textual description?
-In this work, we introduce 3M-Diffusion, a novel multi-modal molecular diffusion method, for generating high-quality, diverse, novel molecular graphs given a textual description.
-Inspried by text-guided image generation using a diffusion model on the latent space (Rombach et al., 2022), 3M-Diffusion is trained to generate molecules from the latent distribution of a molecular graph autoencoder (See Figure 1).
-This enables the diffusion process to focus on the high-level semantics of molecules.
-We find that this approach is particularly well-suited for discrete molecular graph modality because it offloads the challenge of modeling a discrete and diverse distribution to the autoencoder and simplifies the diffusion process by restricting it to the continuous, latent feature space.
-Since graph data consists of both node attributes and graph structure, which differs significantly from the sequential and contextual nature of text data, to align the latent space of the text and molecular graph, 3M-diffusion first pretrains molecular graph and text encoders through contrastive learning on a large dataset of molecule-text pairs, enabling generate high-quality molecular graphs that match the given textual descriptions.
-Our extensive experiments demonstrate that 3M-Diffusion can generate high-quality, novel, and diverse molecular graphs that semantically match the textual description provided.
+We introduce 3M-Diffusion, a novel multi-modal molecular diffusion method, for generating high-quality, diverse, novel molecular structures from a textual description.
+Inspired by text-guided image generation using a diffusion model on the latent space (Rombach et al., 2022), 3M-Diffusion is trained to generate molecular structures from the latent distribution of a molecular graph autoencoder (See Figure 1).
+This enables the diffusion process to focus on the high-level semantics of encodings of molecular structures.
+We find that this approach is particularly well-suited for discrete molecular graph modality because it relegates the challenge of modeling a discrete and diverse distribution to the autoencoder and simplifies the diffusion process by restricting it to the continuous, latent feature space.
+Since graph data consists of both node attributes and graph structure, which differs significantly from the sequential text data, to align the latent representations of textual descriptions and molecular graphs, 3M-diffusion employs contrastive learning on a large data set of pairs of molecular structures and their textual descriptions to pretrain molecular graph encoders and LLM text encoders.
+This enables 3M-diffusion to generate high-quality molecular graphs that match the given textual descriptions.
 Key Contributions.
-The key contributions of this paper are as follows: (i) To the best of our knowledge, 3M-Diffusion offers the first multimodal diffusion approach to generating molecular graphs from their textual descriptions, surpassing the limitations of state-of-the-art (SOTA) methods for this problem.
-(ii) 3M-Diffusion aligns the latent spaces of molecular graphs and textual descriptions to offer a text-molecule aligned latent diffusion model to generate higher-quality, diverse, and novel molecular graphs that match the textual description provided; and (iii) We present results of extensive experiments using four real-world text-based molecular graphs generation benchmarks that show that 3M-Diffusion outperforms SOTA methods for both text-guided and unconditional molecular graph generation.
+The key contributions of this paper are as follows: (i) To the best of our knowledge, 3M-Diffusion offers the first multimodal diffusion approach to generating molecular structures from their textual descriptions, surpassing the limitations of state-of-the-art (SOTA) methods for this problem.
+(ii) 3M-Diffusion aligns the latent spaces of molecular graphs and textual descriptions to offer a text-molecule aligned latent diffusion model to generate higher-quality, diverse, and novel molecular structures that match the textual description provided; and (iii) Results of extensive experiments using four real-world text-based molecular graphs generation benchmarks show that 3M-Diffusion outperforms SOTA methods for both text-guided and unconditional molecular structure generation.
 In particular, 3M-Diffusion achieves 146.27% novelty, 130.04% diversity relative improvement over the SOTA method with maintained semantics in textual prompt on the PCDes dataset.
+Figure 2: Qualitative comparisons to the MolT5-large of generated molecules on ChEBI-20.
+Compared with the SOTA method MolT5-large, our results are more diverse and novel with maintained semantics in textual prompt.
+More results are provided in Appendix D.4
 2 Related Work
-2.1 Molecule Generation
+2.1 Molecular Structure Generation
 Existing molecular graph generation methods can be broadly grouped into two categories: text-based and graph-based models.
 Text-based models typically utilize the SMILES (Weininger, 1988) or SELFIES (Krenn et al., 2022) linear strings to describe each molecule  (Bjerrum & Threlfall, 2017; Gómez-Bombarelli et al., 2018; Kusner et al., 2017; Flam-Shepherd et al., 2022; Fang et al., 2023b; Grisoni, 2023).
-Because a given molecular graph can have several linear string representations and slight changes to the string representation can yield significant changes in molecular graph being described, such linear string encodings are far from ideal for learning generative models for producing molecular graph from textual descriptions, several authors have used deep generative models, including graph variational autoencoders (VAEs) (You et al., 2018; Jin et al., 2018; Xiao et al., 2021; Jin et al., 2020; Kong et al., 2022; Liu et al., 2018; Diamant et al., 2023), normalizing flows (Madhawa et al., 2019; Zang & Wang, 2020; Luo et al., 2021), generative adversarial networks (Guarino et al., 2017; Maziarka et al., 2020), diffusion models (Niu et al., 2020; Vignac et al., 2022; Jo et al., 2022) and autoregressive models (You et al., 2018; Xiao & Wang, 2021; Popova et al., 2019; Goyal et al., 2020) to learn distributions of molecular graphs from large databases of known graphs.
-However, many of these studies focus on generating molecular graphs with specified low-level properties, such as logP (the octanol-water partition coefficient) as opposed to those that match high-level textual descriptions of molecules (Edwards et al., 2022).
-However, practical applications, such as drug design, call for effective approaches to generating diverse and novel molecular graphs that match a given textual description.
-2.2 Text-guided Generation of Molecular Graphs
+Because a given molecular graph can have several linear string representations and small changes to the string representation can result in large changes in molecular graph being described, such linear string encodings are far from ideal for learning generative models for producing molecular graph from textual descriptions.
+Inspired by the success of learning enhanced graph representations using graph based models (Kipf & Welling, 2016; You et al., 2020; Xiao et al., ; Xiao et al., 2024), several authors have employed deep generative models, including graph variational autoencoders (VAEs) (You et al., 2018; Jin et al., 2018; Xiao et al., 2021; Jin et al., 2020; Kong et al., 2022; Liu et al., 2018; Diamant et al., 2023), normalizing flows (Madhawa et al., 2019; Zang & Wang, 2020; Luo et al., 2021), generative adversarial networks (Guarino et al., 2017; Maziarka et al., 2020), diffusion models (Niu et al., 2020; Vignac et al., 2022; Jo et al., 2022) and autoregressive models (You et al., 2018; Xiao & Wang, 2021; Popova et al., 2019; Goyal et al., 2020) to learn distributions of molecular graphs from large databases of known graphs.
+However, many of these studies focus on generating molecular graphs with specific low-level properties, such as logP (the octanol-water partition coefficient) as opposed to those that match high-level textual descriptions of a broader range of molecular properties  (Edwards et al., 2022), e.g., water solubility, chemical activity, etc.
+However, practical applications, such as drug design, call for effective approaches to generating diverse and novel molecular graphs that match a given textual description of such high-level properties.
+2.2 Language-guided Molecule Generation
 Advances in deep learning and language models have inspired a growing body of work on applications of such models to problems in the molecular sciences  (Schwaller et al., 2019; Xiao et al., 2023; Toniato et al., 2021; Xiao et al., 2024; Zhao et al., 2023a; Seidl et al., 2023; Schwaller et al., 2021; Vaucher et al., 2020).
 Advances in text-guided generation of images, videos, and audios  (Radford et al., 2021; Rombach et al., 2022; OpenAI, 2023; Yin et al., 2023), together with the flexibility offered by natural language descriptions of molecular graphs have inspired a growing body of work on text-guided generation of molecular graphs (Christofidellis et al., 2023; Edwards et al., 2022; Fang et al., 2023a; Zeng et al., 2022; Su et al., 2022).
-For example, MolT5 (Edwards et al., 2021, 2022) and ChemT5 (Christofidellis et al., 2023) pre-train on a substantial volume of unlabeled language text and SMILES strings for tasks such as molecule captioning and text-based molecule generation (Christofidellis et al., 2023; Edwards et al., 2022).
+For example, MolT5 (Edwards et al., 2021; 2022) and ChemT5 (Christofidellis et al., 2023) pre-train on a substantial volume of unlabeled language text and SMILES strings for tasks such as molecule captioning and text-based molecule generation (Christofidellis et al., 2023; Edwards et al., 2022).
 However, as noted above, the degenerate nature of linear SMILES string representation of molecular graphs makes it less ideal for generating molecules from their textual descriptions.
-To address this problem, we propose 3M-Diffusion, aligning latent representations of textual descriptions for molecules with the latent representations of the graphs themselves, offering a powerful approach for text-guided molecule generation.
+To address this problem, we propose 3M-Diffusion, a novel method that aligns the latent representation of textual descriptions of molecules with that of the corresponding molecular structures, yielding a powerful approach to text-guided molecular structure generation.
 3 Preliminaries
 3.1 Problem Definition
-In this paper, we consider generative modeling of 2D molecular graphs from scratch.
+We consider generative modeling of 2D molecular graphs.
 Each molecule is represented as a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}=\left\{v_{1},\ldots,v_{|\mathcal{V}|}\right\}$ is the set of $|\mathcal{V}|$ nodes and $\mathcal{E}$ is the set of edges.
 Let $\mathbf{X}=\left[\mathbf{x}_{1},\mathbf{x}_{2},\cdots,\mathbf{x}_{|\mathcal{V}\mid}\right]\in\mathbb{R}^{\left[\mathcal{V}\mid\times D_{x}\right.}$ be the node attribute matrix, where $x_{i}$ is the $D_{x}$-dimensional one-hot encoding feature vector of $v_{i}$, such as atomic type and chirality type.
-Similarly, a tensor $\mathbf{A}\in\mathbb{R}^{|\mathcal{V}|\times|\mathcal{V}|\times b}$ groups the one-hot encoding $\mathbf{e}_{ij}$ of each edge, with each entry representing a distinct edge type (bond types for molecule), while treating the absence of an edge as a specific edge type.
-Given the dataset $\mathcal{D}=\{(\mathcal{G}_{i},\mathcal{T}_{i})_{i=1}^{N}\}$ where molecule $\mathcal{G}$ has specific text descriptions $\mathcal{T}$ characterizing the features of the molecule, our goal is to learn conditional generation models $p_{\theta}(\mathcal{G}\mid\mathcal{T})$ for generating novel, diverse, and valid molecules while simultaneously aligning with the desired features outlined in text descriptions $\mathcal{T}$.
-For brevity, in what follows, we omit the input subscript $i$ unless otherwise stated.
+Similarly, a tensor $\mathbf{A}\in\mathbb{R}^{|\mathcal{V}|\times|\mathcal{V}|\times b}$ groups the one-hot encoding $\mathbf{e}_{ij}$ of each edge, where each entry is a distinct edge type (bond types for molecule), with the absence of an edge encoded explicitly using a designated edge type.
+Given the dataset $\mathcal{D}=\{(\mathcal{G}_{i},\mathcal{T}_{i})_{i=1}^{N}\}$ where molecule $\mathcal{G}$ has specific text descriptions $\mathcal{T}$ characterizing the features of the molecule, our goal is to learn conditional generation models $p_{\theta}(\mathcal{G}\mid\mathcal{T})$ for generating novel, diverse, and valid molecules while also aligning with the desired features outlined in text descriptions $\mathcal{T}$.
+For brevity, in what follows, we omit the input subscript $i$ when it is clear from context.
 3.2 Diffusion Models
@@
 After training, we can draw samples with $\hat{\mathbf{z}}_{\theta}$ by the iterative ancestral sampling (Ho et al., 2020):
-$\mathbf{z}_{t-1}$ $=\tfrac{\sqrt{\alpha_{t-1}}\left(1-\alpha_{t\mid t-1}\right)}{1-\alpha_{t}}\hat{\mathbf{z}}_{\theta}(\mathbf{x}_{t},t)$ (2)
-$+\tfrac{\sqrt{\alpha_{t\mid t-1}}\left(1-\alpha_{t-1}\right)}{1-\alpha_{t}}\mathbf{z}_{t}+\sigma(t-1,t){\bm{\epsilon}},$
+$\mathbf{z}_{t-1}$ $=\tfrac{\sqrt{\alpha_{t-1}}\left(1-\alpha_{t\mid t-1}\right)}{1-\alpha_{t}}\hat{\mathbf{z}}_{\theta}(\mathbf{x}_{t},t)+\tfrac{\sqrt{\alpha_{t\mid t-1}}\left(1-\alpha_{t-1}\right)}{1-\alpha_{t}}\mathbf{z}_{t}+\sigma(t-1,t){\bm{\epsilon}},$ (2)
 where $\alpha_{t\mid t-1}$ represents $\alpha_{t}/\alpha_{t-1}$ and $\sigma(t-1,t)=(1-\alpha_{t-1})\left(1-\alpha_{t\mid t-1}\right)/(1-\alpha_{t})$.
@@
 4 Latent Multi-Modal Diffusion for Molecules
-Our 3M-Diffusion in Figure 1 aims to learn a probabilistic mapping from the latent space of textual descriptions to a latent space of molecular graphs for text-guided generation of molecular graphs.
+3M-Diffusion (See Figure 1) aims to learn a probabilistic mapping from the latent space of textual descriptions to a latent space of molecular graphs for text-guided generation of molecular graphs.
 However, direct mapping between these two latent representations using a diffusion model faces a key hurdle, namely, large mismatch between the latent spaces of textual descriptions and of molecular graphs.
-To overcome this hurdle, 3M-Diffusion adopts a two-part approach using a text-molecule aligned variational autoencoder and a multi-modal molecule latent diffusion.
-The former learns to align the representation of molecular graphs with that of their textual descriptions using contrastive learning, followed by a decoder that maps the latent representation back to the corresponding molecular graphs.
-The latter works with the text-aligned latent representation of molecular graphs to learn a conditional generative model, that maps textual descriptions to latent representations of molecular graphs.
-As we will see later, 3M-Diffusion is able to produce a diverse collection of novel molecular graphs that match the given textual descriptions.
+To overcome this hurdle, 3M-Diffusion adopts a two-step approach.
+The first step uses contrastive learning to produce text-molecule aligned variational autoencoder that align the representation of molecular graphs with that of their textual descriptions, and a decoder that maps the latent representation back to the corresponding molecular graphs.
+The second step uses the text-aligned latent representation of molecular graphs to learn a conditional generative model, that maps textual descriptions to latent representations of molecular graphs.
+As we will see later, 3M-Diffusion can produce diverse novel molecular graphs that match the given textual descriptions.
 4.1 Text-Molecule Aligned Variational Autoencoder
-Our text-molecule aligned variational autonecoder comprises of three parts: the molecular graph encoder $E_{g}$, text encoder $E_{t}$ and molecular graph decoder $D$.
+The proposed text-molecule aligned variational autonecoder comprises of three parts: the molecular graph encoder $E_{g}$, LLM encoder $E_{t}$ and molecular graph decoder $D$.
 To bridge the representation gap between molecular graphs and their textual descriptions, the text and molecular graph encoders are trained on a large-scale dataset of molecule-text pairs.
-These encoders learn a well-aligned shared text-molecule representation, enriching the semantics of the molecular graph representation by aligning it the latent representation of textual descriptins of molecules through contrastive learning.
-These three parts are described in more detail below:
+These encoders use contrastive learning to construct well-aligned text-molecular structure encodings that align the latent representation of molecular structures by aligning it with with that of their textual descriptions.
+The details of this process are given below.
 Molecular Graph Encoder aims to map each molecular graph $\mathcal{G}$ with the adjacency tensor $\mathbf{A}$ and node features matrix $\mathbf{X}$ into a lower-dimensional latent space $\mathcal{Z}$.
@@
 Then, the latent graph representation $\mathbf{z}$ is sampled from a Gaussian distribution as $\mathbf{z}\sim q\left(\mathbf{z}\mid\mathbf{A},\mathbf{X}\right)\simeq\mathcal{N}\left(\bm{\mu}_{g},\bm{\sigma}_{g}\right)$.
-This graph encoding process can also be expressed as $\mathbf{z}=E_{g}(\mathcal{G})$ signifying the outcome of the encoding operation within the input context of $\mathcal{G}$.
-Text Encoder maps the textual descriptions of molecular graphs into a latent space, that provides input to a conditional generative model of molecular graphs.
-This encoded information becomes instrumental in generating molecules that match a given textual description.
-To inject potentially useful scientific knowledge from the literature into $E_{t}$ we initialize it with Sci-BERT (Beltagy et al., 2019), an encoder-only transformer pretrained on scientific publications.
+This graph encoding process can be expressed as $\mathbf{z}=E_{g}(\mathcal{G})$ signifying the outcome of the encoding operation in the input context of $\mathcal{G}$.
+LLM Encoder maps the textual descriptions of molecular structures into a latent space that provides input to a conditional generative model of molecular graphs.
+This encoded information is instrumental in generating molecules that match a given textual description.
+To inject potentially useful scientific knowledge from the literature into $E_{t}$ we initialize it with encoder-based LLM, i.e., Sci-BERT (Beltagy et al., 2019), which is pretrained on the text of scientific publications.
 Aligning the representation of molecular graphs with that of their textual descriptions is crucial for establishing a shared representation for effective text-guided molecular graph generation.
-Graph data consists of both node attributes and graph structure, which differs significantly from the sequential and contextual nature of text data.
-We aim to align the molecule representation from the molecule graph encoder $E_{g}$ with text representation from the text encoder $E_{t}$ by pretraining.
+Graph data encodes both node attributes and graph structure which differs significantly from the sequential and contextual nature of text data.
+We aim to align the molecular representation from the molecule graph encoder $E_{g}$ with text representation from the LLM encoder $E_{t}$ using contrastive learning.
 Consider a molecular graph $\mathcal{G}$ and its textual description $\mathcal{T}$.
@@
 $\cos(\cdot,\cdot)$ denotes cosine similarity and $\tau$ is the temperature hyperparameter.
-The encoders are pretrained on molecule-text (300K) pairs from PubChem collected by  Liu et al.
-(2023) and robust enough to capture a well-aligned molecule-text space, which overcomes the challenges posed by the distribution gap and produce more consistent and high-quality results in cross-modal molecule generation.
-Molecular Graph Decoder learns to use latent representation $\mathbf{z}$ to decode the molecular graph.
-In this work, we employ the HierVAE (Jin et al., 2020) decoder, as it works empirically well in our final model.
-Note that our approach remains flexible and can be easily applied to other graph decoders (Li et al., 2018; Kong et al., 2022; Grover et al., 2019).
-Specifically, it generates a molecule $\mathcal{G}$ using a graph representation $\mathbf{z}$ from the encoder.
-Given the molecular graph $\mathcal{G}$, we minimize the negative Evidence Lower Bound (ELBO) (Jin et al., 2020) to train both encoder and decoder:
+The encoders are pretrained on molecule-text (300K) pairs from PubChem  Liu et al.
+(2023).
+As we will see, this process yields a well-aligned molecule-text space that can support the generation of diverse and novel molecular structures from their textual descriptions.
+Molecular Graph Decoder learns to use latent representation $\mathbf{z}$ to generate the corresponding molecular structure.
+In this work, we employ the HierVAE (Jin et al., 2020) decoder.
+Note however that our approach can be combined with other graph decoders (Li et al., 2018; Kong et al., 2022; Grover et al., 2019).
+Specifically, the molecular graph decoder generates a molecular structure $\mathcal{G}$ using a graph representation $\mathbf{z}$ from the encoder.
+Given the molecular graph $\mathcal{G}$, we minimize the negative Evidence Lower Bound (ELBO) (Jin et al., 2020) to train both the encoder and the decoder:
 $\mathcal{L}_{\text{elbo}}=-\mathbb{E}_{q(\mathbf{z}\mid\mathbf{A},\mathbf{X})}[p(\mathcal{G}\mid\hat{\mathcal{G}})]+\alpha\mathrm{KL}[q(\mathbf{z}\mid\mathcal{G})\|p(\mathbf{z})].$ (5)
@@
 Here, $p(\mathbf{z})=\mathcal{N}\left(\mathbf{z}\mid\mathbf{0},\mathbf{I}\right)$ is the isotropic multivariate Gaussian prior, where $\mathbf{I}$ is the identity matrix.
-We optimize this first expectation term concurrently through reparameterization trick (Kingma & Welling, 2013).
+We optimize this first expectation term concurrently using the reparameterization trick (Kingma & Welling, 2013).
 4.2 Multi-modal Molecule Latent Diffusion
-We then introduce latent diffusion for conditional molecular graph generation, a method that leverages the aligned molecule-text latent space of encoder-decoder discussed above, enabling us preserve the crucial molecular properties and aligning the textual description of $\mathcal{G}$.
-Specifically, we learn a better probabilistic mapping from the text to the aligned moelcule latent space, and thereby generate a higher-quality and more diverse molecules with more semantic consistency conforming to the conditional text inputs.
-Specifically, the denoising network $\hat{\mathbf{z}}_{\theta}$ focuses on generating latent graph representation $\mathbf{z}$ conditioned on the text representation $\mathbf{c}$ from the text encoder $E_{t}$.
-The objective is:
-$\small\mathcal{L}_{\text{diff}}=\mathbb{E}_{t,\mathbf{x}_{0},\epsilon}\left[\lambda_{t}\left\|\hat{\mathbf{z}}_{\theta}\left(\sqrt{\alpha_{t}}\mathbf{z}+\sqrt{1-\alpha_{t}}{\bm{\epsilon}},t,\mathbf{c}\right)-\mathbf{z}\right\|_{2}^{2}\right],$ (6)
+We next introduce latent diffusion for conditional molecular graph generation, a method that leverages the aligned molecule-text latent space of encoder-decoder described above, enabling us to preserve the crucial molecular properties and aligning the textual description of $\mathcal{G}$.
+Specifically, we learn a probabilistic mapping from the text to the text aligned molecular structure latent space, and thus generate molecular structures that match the conditional text inputs.
+The denoising network $\hat{\mathbf{z}}_{\theta}$ focuses on generating latent graph representation $\mathbf{z}$ conditioned on the text representation $\mathbf{c}$ from the LLM encoder $E_{t}$.
+The objective to be optimized is:
+$\mathcal{L}_{\text{diff}}=\mathbb{E}_{t,\mathbf{x}_{0},\epsilon}\left[\lambda_{t}\left\|\hat{\mathbf{z}}_{\theta}\left(\sqrt{\alpha_{t}}\mathbf{z}+\sqrt{1-\alpha_{t}}{\bm{\epsilon}},t,\mathbf{c}\right)-\mathbf{z}\right\|_{2}^{2}\right],$ (6)
 where $\lambda_{t}$ are time-dependent weights and we use MLPs for the network $\hat{\mathbf{z}}_{\theta}$.
@@
 $\mathbf{c}$ is the text representation from the fixed encoder $E_{t}$ pretrained by Equation (4).
-In addition, we improve sample quality through classifier-free guidance (Ho & Salimans, 2022), jointly training an unconditional network, $\hat{\mathbf{z}}_{\theta}(\mathbf{z}_{t},t)$, and a conditional network, $\hat{\mathbf{z}}_{\theta}\left(\mathbf{z}_{t},t,\mathbf{c}\right)$.
+Additionally, to improve sample quality we take advantage of classifier-free guidance (Ho & Salimans, 2022), to jointly train an unconditional network, $\hat{\mathbf{z}}_{\theta}(\mathbf{z}_{t},t)$, and a conditional network, $\hat{\mathbf{z}}_{\theta}\left(\mathbf{z}_{t},t,\mathbf{c}\right)$.
 Conditioning information is randomly dropped with a probability of $p=0.1$ during training.
-When conditioning information is dropped, we replace the embedded source text with a learnable embedding.
+When conditioning information is dropped, we replace the embedded source text with its embedding.
 4.3 Training and Inference
 Training.
-We now describe the training procedure for 3M-Diffusion.
-While the learning objectives for graph encoder-decoder and multi-modal latent diffusion have already been specified by Equations (5) and (6), it is unclear whether the two components should be trained sequentially
+We proceed to describe the training procedure for 3M-Diffusion.
+While the learning objectives for graph encoder-decoder and multi-modal latent diffusion have already been specified by Equations (5) and (6), it is unclear whether the two components should be trained sequentially.
 Previous research on latent diffusion models for image generation indicates that a two-stage training strategy often results in superior performance (Rombach et al., 2022).
 Hence, we adopt a similar two-stage strategy.
-Specifically, our first stage is to train the encoder-decoder using Equation (5), and the second stage is to train multi-modal molecule latent diffusion with the fixed encoder-decoder from stage one.
+The first stage consists of training an encoder-decoder pair using Equation (5), and the second stage consists of training a multi-modal molecule latent diffusion with the fixed encoder-decoder from the first stage.
 Inference. During inference, we use a weight $w$ in computing the prediction as follows (Ho & Salimans, 2022):
@@
 The new sampling equation is:
-$\mathbf{z}_{t-1}$ $=\tfrac{\sqrt{\alpha_{t-1}}\left(1-\alpha_{t\mid t-1}\right)}{1-\alpha_{t}}\tilde{\mathbf{z}}_{t}$ (8)
-$+\tfrac{\sqrt{\alpha_{t\mid t-1}}\left(1-\alpha_{t-1}\right)}{1-\alpha_{t}}\mathbf{z}_{t}+\sigma(t-1,t){\bm{\epsilon}},$
+$\mathbf{z}_{t-1}$ $=\tfrac{\sqrt{\alpha_{t-1}}\left(1-\alpha_{t\mid t-1}\right)}{1-\alpha_{t}}\tilde{\mathbf{z}}_{t}+\tfrac{\sqrt{\alpha_{t\mid t-1}}\left(1-\alpha_{t-1}\right)}{1-\alpha_{t}}\mathbf{z}_{t}+\sigma(t-1,t){\bm{\epsilon}},$ (8)
 where the sampling chain is initialized from Gaussian prior $\mathbf{z}_{T}\sim p\left(z_{T}\right)=\mathcal{N}\left(\mathbf{z}_{T};\mathbf{0},\bm{I}\right)$.
-The detailed training and inference algorithms of 3M-diffusion are in Appendix B.
-Table 1: Quantitative comparison of conditional generation on the ChEBI-20 and PubChem.
+The detailed training and inference algorithms of 3M-diffusion are provided in Appendix B.
+Table 1: Quantitative comparison of conditional generation on the PCDes and MoMu.
 3M-Diffusion outperforms other SOTA methods in terms of novelty, diversity, and validity metrics by a large margin, while maintaining a good similarity metric.
@@
 ChemT5-base 85.01 25.55 14.08 92.93 23.40 97.65 30.07 87.61
+Mol-Instruction 60.86 35.60 24.57 79.19 14.89 97.52 30.17 68.32
 3M-Diffusion 81.57 63.66 32.39 100.0 24.62 98.16 37.65 100.0
@@
 ChemT5-base 95.48 15.12 13.91 97.15 89.42 22.40 13.98 92.43
+Mol-Instruction 65.75 32.01 26.50 77.91 23.40 37.37 27.97 71.10
 3M-Diffusion 87.09 55.36 34.03 100 87.05 64.41 33.44 100
-5 Experiment
-5.1 Experimental Settings
-Table 3: Statistics of all datasets.
-Dataset #Training #Validation #Test
-ChEBI-20 15,409 1,971 1,965
-PubChem 6,912 571 1,162
-PCDes 7,474 1,051 2,136
-MoMu 7,474 1,051 4,554
+5 Experiments and Results
+5.1 Experimental Setup
 Datasets.
-In our experiments, we use PubChem (Liu et al., 2023), ChEBI-20 (Edwards et al., 2021), PCDes (Zeng et al., 2022) and Momu (Su et al., 2022) datasets.
-Following the evaluation methodology adopted with datasets (ZINC250K (Irwin et al., 2012) and QM9 (Blum & Reymond, 2009; Rupp et al., 2012)) used in previous studies of automated generation of molecular structures, we limit our training, validation, and test data to molecules with fewer than 30 atoms.
+We use PubChem (Liu et al., 2023), ChEBI-20 (Edwards et al., 2021), PCDes (Zeng et al., 2022) and Momu (Su et al., 2022) datasets.
+Following the evaluation methodology adopted with datasets (ZINC250K (Irwin et al., 2012) and QM9 (Blum & Reymond, 2009; Rupp et al., 2012)) used in previous studies of molecular structure generation, we limit our training, validation, and test data to molecules with fewer than 30 atoms.
 Statistics of datasets are shown in Table 5 and detailed descriptions are given in the Appendix C.1.
@@
 We evaluate the model’s performance on two tasks: text-guided and unconditional molecule generation.
-Following previous works on text-guided molecule generation, we compare our 3M-Diffusion the SOTA baselines (Edwards et al., 2022; Christofidellis et al., 2023).
+Following previous works on text-guided molecule generation, we compare our 3M-Diffusion the SOTA baselines (Edwards et al., 2022; Christofidellis et al., 2023; Fang et al., 2023a).
 Text-guided molecule generation measures the model’s capacity to generate chemically valid, structurally diverse, novel molecules that match the textual description provided.
-Based on previous work (Chen et al., 2021; Fu et al., 2021; Wang et al., 2022), we adopt the following metrics for this task: Similarity, Novelty, Diversity, Validity.
-We call a predicted molecular graph, specifically one that is chemically valid ($g^{\prime}$), qualified with the ground truth graph $g$ if $f(g,g^{\prime})>0.5$, where $f$ measures the MACCS (Durant et al., 2002) cosine similarity between the two.
+As in previous work (Chen et al., 2021; Fu et al., 2021; Wang et al., 2022), we adopt the following metrics for this task: Similarity, Novelty, Diversity, Validity.
+We call a predicted molecular structure, specifically one that is chemically valid ($g^{\prime}$), qualified with the ground truth graph $g$ if $f(g,g^{\prime})>0.5$, where $f$ measures the MACCS (Durant et al., 2002) cosine similarity between the two.
 If the $f(g,g^{\prime})$ similarity between $g^{\prime}$ and $g$ is smaller than a threshold (0.8 in this paper), we call this predicted molecule $g$ novel.
@@
 (iv) Validity: the percentage of generated molecules that are chemically valid out of all molecules.
-We report the results for unconditional molecule generation tasks using the GuacaMol benchmarks (Brown et al., 2019) to assess the model’s capability to learn molecular data distribution and generate realistic, diverse molecules.
+We report the results for unconditional molecule generation tasks using the GuacaMol benchmarks (Brown et al., 2019) to assess the model’s ability to learn a generative model of molecular structures that can be used to generate realistic, diverse molecules.
 Specifically, Uniqueness gauges the ratio of unique molecules among the generated ones.
@@
 Each metric is normalized to a range of 0 to 1, with a higher value indicating better performance.
+Baselines.
+For text guided molecule generation, we compare 3M-Diffusion with the following baselines: MolT5 (Edwards et al., 2022), ChemT5 (Christofidellis et al., 2023) and Mol-Instruction (Fang et al., 2023a).
+We consider seven representative baselines for unconditional molecule generation based on Variational Autoencoder, including CharRNN (Segler et al., 2018), VAE (Kingma & Welling, 2013), AAE (Makhzani et al., 2015), LatentGAN (Prykhodko et al., 2019) BwR (Diamant et al., 2023), HierVAE (Jin et al., 2020) and PS-VAE (Kong et al., 2022).
+The details of the baselines are given in Appendix C.1.
+Setup.
+We randomly initialize the model parameters and use the Adam optimizer to train the model.
+HierVAE is used as our variational autoencoder for all of the following experiments.
+For baselines of unconditional generation that did not report results in these datasets, we reproduce the results using the official code made available by the authors.
+To ensure a fair comparison across all methods for unconditional generation, we select the best hyperparameter configuration solely based on the loss on the training set.
+We evaluate the performance of baselines for text-guided molecule generation using the pretrained parameters made available by the authors of the respective studies.
+5.2 Text-guided Molecule Generation
+Figure 3: Inference time comparison for conditional molecule generation on ChEBI-20.
+Table 1 and 2 report Similarity, Novelty, Diversity, and Validity for text-guided molecule generation on four datasets.
+Each metric is computed as an average over 5 structures generated from a given textual description.
+We see that 3M-Diffusion is competitive with SOTA baselines with respect to Similarity and outperforms them substantially with respect to Diversity and Novelty.
+Specifically, 3M-Diffusion achieves 146.27% (Novelty of PCDes), 130.04% (Diversity of PCDes) relative improvement over the second-best SOTA baseline while matching the textual prompt.
+We conclude that while the transformer based models that are pretrained on a large dataset excel in producing molecular structures that match the textual descriptions provided, 3M-Diffusion outperforms them with respect to the diversity and novelty of the molecular structures without sacrificing the match between the textual descriptions provided as input and the molecular structures generated as output.
+Also, as Figure  3 shows, 3M-Diffusion enjoys a substantial advantage in terms of its speed over transformer-based methods (See Appendix D.1 for details).
 Table 3: Quantitative comparison of unconditional generation.
@@
 # Methods Uniq (%) Novelty (%) KL Div (%) FCD (%) Validity (%) Uniq (%) Novelty (%) KL Div (%) FCD (%) Validity (%)
+CharRNN 72.46 11.57 95.21 75.95 98.21 63.28 23.47 90.72 76.02 94.09
+VAE 57.57 47.88 95.47 74.19 63.84 44.45 42.47 91.67 55.56 94.10
+AAE 1.23 1.23 38.47 0.06 1.35 2.94 3.21 39.33 0.08 1.97
+LatentGAN 66.93 57.52 94.38 76.65 73.02 52.00 50.36 91.38 57.38 53.62
 BwR 22.09 21.97 50.59 0.26 22.66 82.35 82.34 45.53 0.11 87.73
@@
 3M-Diffusion 83.04 70.80 96.29 77.83 100.0 85.42 81.20 92.67 58.27 100.0
-Baselines.
-For text guided molecule generation, we compare 3M-Diffusion with the following baselines: MolT5 (Edwards et al., 2022) and ChemT5 (Christofidellis et al., 2023).
-We consider three representative baselines for unconditional molecule generation based on Variational Autoencoder, including BwR (Diamant et al., 2023), HierVAE (Jin et al., 2020) and PS-VAE (Kong et al., 2022).
-The descriptions and details of the baselines are given in Appendix C.1.
-Setup.
-We randomly initialize the model parameters and use the Adam optimizer to train the model.
-HierVAE is used as our variational autoencoder for all of the following experiments.
-For baselines of unconditional generation that did not report results in these datasets, we reproduce the results using the official code of the authors.
-To ensure a fair comparison across all methods for unconditional generation, we select the best hyperparameter configuration solely based on the loss on the training set.
-We evaluate the performance of baselines for text-guided molecule generation using the released pre-trained parameters.
-Figure 3: Molecules generated unconditionally by 3M-diffusion from the prior distribution $\mathcal{N}(\mathbf{0},\bm{I})$.
-5.2 Text-guided Molecule Generation
-Table 1 and 2 report Similarity, Novelty, Diversity, and Validity for text-guided molecule generation on four datasets.
-All of these metrics are computed as averages, with each metric being calculated based on 5 samples generated from each textual description.
-Our model demonstrates competitive results in Similarity and notably higher values in the Diversity and Novelty.
-Specifically, 3M-Diffusion achieves 146.27% (Novelty of PCDes), 130.04% (Diversity of PCDes) relative improvement over the second-best method with maintaining semantics in the textual prompt.
-This suggests that transformer-based models, pre-trained on a large dataset, excel in predicting molecules based on textual descriptions.
-However, our model can generate a diverse and novel collection of molecular graphs that match the textual descriptions provided as input.
-Also, as Figure  4 shows, our model can generate molecules faster than transformer-based methods.
 5.3 Unconditional Molecule Generation
 Table 3 shows the results of unconditional molecule generation on PubChem and ChEBI-20.
-3M-Diffusion consistently outperforms or achieves competitive results across all four metrics on both datasets, indicating its ability to generate realistic molecules while avoiding overfitting the training data.
+3M-Diffusion consistently outperforms or matches the performance of SOTA baselines across all four metrics on both datasets, suggesting that it is able to generate realistic, diverse, and novel molecules without overfitting the training data.
 Furthermore, 3M-Diffusion consistently outperforms HierVAE when both methods use the same decoder.
-This underscores the clear advantage of 3M-Diffusion in generating a diverse and novel assortment of molecular graphs.
-As shown in Figure 3, our sampled molecules present rich variety and structural complexity.
-We have included more sampled molecules from the prior distribution in Appendix D.4.
-Figure 4: The average inference time comparison of molecule conditional generation on ChEBI-20 dataset.
-5.4 Ablation Study
-We conduct ablation studies to assess the impact of different model designs, including joint training of the diffusion model and Variational Autoencoder (denoted as ”Without two stage training”) and training the model without aligning the representation space of graphs and text using $\mathcal{L}_{\text{con}}$ in Equation (4) (denoted as ”Without representation alignment”).
-Additionally, we perform experiments without using either of these configurations, labeled as ”Without two stage training & Without representation alignment” in Table 5.
-Initially, training the aligned encoder-decoder alongside Multi-modal Molecule Latent Diffusion jointly yielded comparable or superior performance in terms of diversity and novelty.
-However, there was a noticeable deficiency in its ability to learn align its representation of molecular graphs with that of their textual descriptions.
-Training 3M-Diffusion without the aligned encoder-decoder results in substantial drop in performance relative to 3M-Diffusion.
-This result underscores the importance of aligning the representation of textual descriptions of molecules with that of molecular graphs.
-Without using both design, Similarity suffers a substantial decline to 10.24, rendering it unsuitable for real-world applications.
+This underscores the clear advantage of 3M-Diffusion in generating a diverse and novel molecular graphs (See Figure 10 for 50 samples from the learned distribution).
+5.4 Ablation Studies
+We conduct ablation studies to assess the impact of different model designs, by joint training of the diffusion model and Variational Autoencoder (”without two-stage training”) and training the model without aligning the representation space of graphs and text using $\mathcal{L}_{\text{con}}$ in Equation (4) ( ”without representation alignment”).
+Additionally, we perform experiments without both two-stage training and representation alignment (”without two-stage training & representation alignment”) in Table 4.
+We find that joint training the aligned encoder-decoder alongside Multi-modal Molecule Latent Diffusion achieves comparable or superior performance relative to 3M-Diffusion in terms of diversity and novelty, but at the expense of reduced quality of match between the molecular structures produced and the textual descriptions provided.
+We find that training the model without the aligned encoder-decoder results in substantial drop in performance relative to 3M-Diffusion, underscoring the importance of aligning the representation of textual descriptions of molecules with that of molecular structures.
+Ablation of both two-stage training and representation alignment components leads to a substantial drop in performance (Similarity of 10.24), rendering the resulting model unsuitable for real-world applications.
 The full 3M-Diffusion model (last row) achieves the best performance with respect to all of the metrics.
-This confirms the suitability of the two-step training approach for generating diverse, novel set of molecules based on their textual descriptions.
-Table 5: Results of ablation study with different model designs of 3M-diffusion for the Pubchem dataset.
+This confirms the importance of two-step training and representation alignment in training a model that can generate diverse, novel set of molecular structures from their textual descriptions.
+Table 4: Results of ablation study with different model designs of 3M-diffusion for Pubchem.
 # Methods Similarity (%) Novelty (%) Diversity (%) Validity (%)
@@
 5.5 Qualitative Analysis and Case Study
-Figure 5: Molecules generated conditionally on input text by 3M-Diffusion on the ChEBI-20 dataset.
 Property Conditional Molecule Generation.
@@
 logP assesses a molecule’s solubility, while QED evaluates its drug likeness (Bickerton et al., 2012).
-Previous methods have explored ways to optimize their models further in order to generate molecules with desired properties.
-However, in the case of our model, a different approach is taken.
-Our model utilizes designed prompts to realize the conditional molecule generation process, as illustrated in Figure 2 for assessing molecule solubility.
-In particular, we generated 10 samples for each prompt and subsequently select the top 5 samples based on the desired property values.
+Previous methods have explored ways to optimize such models further in order to generate molecules with desired properties.
+In contrast, 3M-Diffusion utilizes textual prompts for conditional molecule generation (See Figure 2).
+In particular, we generated 10 samples for each prompt and subsequently select the top 5 samples based on the value of the desired property.
 For MolT5-large, we adjust the temperature of the transformer in an attempt to encourage diversity while ensuring that invalid samples are controlled during the sampling process.
 We applied the same methodology used in our previous experiments to maintain consistency across the comparative analysis, as shown in Figure 2.
-Compared with the MolT5-large model, we find that 3M-Diffusion excels in generating a diverse range of molecules while achieving improved desired properties, as evidenced by higher logP values Figure 2.
-This underscores our model’s capacity to produce molecules that are not only novel and diverse but also match the textual description provided.
+When compared to the MolT5-large model, we find that 3M-Diffusion excels in generating a diverse range of molecules that also score better with respect to the desired properties, as evidenced by higher logP values Figure 2.
+This underscores our 3M-Diffusion’s ability to produce molecules that are not only novel and diverse but also match the textual description provided.
 This unique approach distinguishes 3M-Diffusion from prior methods for molecular generation, presenting a fresh and promising perspective on molecular design.
-For additional experimental details and insights regarding prompts related to solubility and drug likeness, please refer to the Appendix D.1.
+Additional experimental details and insights regarding prompts related to solubility and drug likeness are provided in Appendix D.2.
+Figure 4: Molecules generated conditionally on input text by 3M-Diffusion on ChEBI-20.
 Qualitative Comparison.
-In Figure 2, we introduce more specific prompts related to molecular categories, such as ”The molecule is a member of the class of flavones.” These prompts bias the model, in favor of generating useful drug candidates.
+In Figure 2, we introduce more specific prompts related to molecular categories, such as ”The molecule is a member of the class of flavones.” These prompts bias the model in favor of generating useful drug candidates.
 Specifically, we choose a reference molecule from the training dataset that matches the given textual prompt.
 Subsequently, we generate ten additional molecules and select the top five molecules that exhibit the highest structural similarity (calculated by Rascal algorithm (Raymond et al., 2002)) to the reference molecule.
-These selected molecules are shown in Figure 2.Comparisong of our model with MolT5-large shows that our model generates more diverse molecules that exhibit a higher degree of structural similarity to the reference molecule.
-Simultaneously, it also produces novel molecular graphs.
-We also include more prompt examples in Appendix D.2.
+These selected molecules are shown in Figure 2.Comparison of 3M-Diffusion with MolT5-large shows that 3M-Diffusion generates more diverse molecules that exhibit a higher degree of structural similarity with the reference molecule while producing novel molecular structures.
+Additional examples of prompts are provided in Appendix D.3.
 Case Study.
-In Figure 5, we choose a set of examples from the ChEBI-20 dataset to visually illustrate the structural similarity between the molecules generated by 3M-Diffusion and the ground truth molecules corresponding to the given text descriptions.
-Specifically, we show the largest substructure that appears in generated and ground truth molecules by RDKit.
-3M-Diffusion demonstrates the capability to generate molecules that share a common substructure, as illustrated in Figure 5.
-Simultaneously, it has the flexibility to explore novel components or structures, thereby achieving a balance between generating similar structures and introducing novel molecules.
-This observation further confirms the ability of our model in generating a diverse collection of novel molecular structures that match the textual description provided.
-We show more results in Appendix D.3.
+In Figure 4, we choose a set of examples from the ChEBI-20 dataset to visually illustrate the structural similarity between the molecules generated by 3M-Diffusion and the ground truth structures corresponding to the given textual descriptions.
+Specifically, we show the largest substructure shared by the generated and ground truth molecules (using RDkit (Landrum, 2010)).
+3M-Diffusion demonstrates the capability to generate molecules that include the relevant substructure, as illustrated in Figure 4.
+Simultaneously, it displays the flexibility to generate molecules with novel substructures, thereby striking a balance between generating structures that are similar to those represented in the training data and suggesting novel structures that match the textual description provided.
+We include additional results in Appendix D.4.
 6 Conclusion
 Generating molecules with desired properties is a critical task with broad applications in drug discovery and materials design.
-These applications call for methods that generate diverse, and ideally novel, molecules with the desired properties.
+Such applications call for methods that produce diverse, and ideally novel, molecules with the desired properties.
 We have introduced 3M-Diffusion, a novel multi-modal molecular graph generation method, to address this challenge.
@@
 3M-Diffusion learns a probabilistic mapping from the text space to the latent molecular graph space using the latent diffusion model.
-The results of our extensive experiments demonstrate that 3M-Diffusion can generate high-quality, novel and diverse molecular graphs that semantically match the textual description provided.
-Ablation studies underscore the importance of aligning the latent representation of texts with that of molecular graphs to the performance of 3M-Diffusion.
-Work in progress is aimed at text-guided molecular graph generation tasks, e.g., text-guided generation of biological macromolecules, new drugs and new materials, and experimental validation of the resulting graphs through their synthesis and characterization in the lab.
+The results of extensive experiments show that 3M-Diffusion can generate high-quality, novel and diverse molecular structures that semantically match the textual description provided.
+Ablation studies underscore the critical role of aligning the latent representation of text with that of molecular graphs in 3M-Diffusion.
+Acknowledgments
+We thank the anonymous reviewers and Prof.
+Dane Morgan of University of Wisconsin-Madison for suggestions that have helped us improve the paper.
+This work was supported in part by grants from the National Science Foundation (2020243, 2226025) to Vasant Honavar and the National Center for Advancing Translational Sciences, and the National Institutes of Health (UL1 TR002014).
+Impact Statements
+This paper advances machine learning approaches to text-guided generation of molecular structures, with potential application in drug design and material discovery.
+The resulting molecules, if successfully synthesized and experimentally characterized in the lab, could provide new drugs for combating diseases and for new materials for wide ranging applications.
 Appendix A Diffusion Models
@@
 With a trained denoising network, we define the mean and variance function in Equation (10) following the previous work (Ho et al., 2020) as:
-$\bm{\mu}_{\theta}\left(\mathbf{z}_{t},t-1,t\right)=\frac{\sqrt{\alpha_{t-1}}\left(1-\alpha_{t\mid t-1}\right)}{1-\alpha_{t}}\hat{\mathbf{z}}_{\theta}(\mathbf{z}_{t},t)+\frac{\sqrt{\alpha_{t\mid t-1}}\left(1-\alpha_{t-1}\right)}{1-\alpha_{t}}\mathbf{z}_{t},\sigma^{2}(t-1,t)=\frac{\left(1-\alpha_{t-1}\right)\left(1-\alpha_{t\mid t-1}\right)}{1-\alpha_{t}},$ (13)
+$\bm{\mu}_{\theta}\left(\mathbf{z}_{t},t-1,t\right)$ $=\frac{\sqrt{\alpha_{t-1}}\left(1-\alpha_{t\mid t-1}\right)}{1-\alpha_{t}}\hat{\mathbf{z}}_{\theta}(\mathbf{z}_{t},t)+\frac{\sqrt{\alpha_{t\mid t-1}}\left(1-\alpha_{t-1}\right)}{1-\alpha_{t}}\mathbf{z}_{t},$ (13)
+$\sigma^{2}(t-1,t)$ $=\frac{\left(1-\alpha_{t-1}\right)\left(1-\alpha_{t\mid t-1}\right)}{1-\alpha_{t}},$
 where $\lambda_{t}$ is the time-dependent weight and $\alpha_{t\mid t-1}=\alpha_{t}/\alpha_{t-1}$.
@@
 The complete training algorithm is depicted in Algorithm 1.
-Firstly, the process begins by training the graph and text encoders using contrastive learning.
+Firstly, the process begins by training the graph and LLM encoders using contrastive learning.
 Following that, the Variational Autoencoder is trained to facilitate the reconstruction of molecule graphs.
@@
 1:  Input: Pairs of Molecule Graph data $\mathcal{G}=(\mathcal{V},\mathcal{E})$ and Textual Description $\mathcal{T}$
-2:  Initial: Graph Encoder network $E_{g}$, Text Encoder network $E_{t}$ and Decoder network $D$, denoising network $\hat{\mathbf{z}}_{\theta}$
+2:  Initial: Graph Encoder network $E_{g}$, LLM Encoder network $E_{t}$ and Decoder network $D$, denoising network $\hat{\mathbf{z}}_{\theta}$
 3:  Pretraining Stage: Aligned Text and Graph Encoder Training
@@
 16:  Second Stage: Multi-modal Molecule Latent Diffusion Training
-17:  Fix Graph Encoder $E_{g}$ from Stage 1 and Text Encoder $E_{t}$ from Pretraining Stage
+17:  Fix Graph Encoder $E_{g}$ from Stage 1 and LLM Encoder $E_{t}$ from Pretraining Stage
 18:  while $\mathbf{z}_{\theta}$ have not converged do
@@
 C.1 Datasets Details and Statistics
+Table 5: Statistics of all datasets.
+Dataset #Training #Validation #Test
+ChEBI-20 15,409 1,971 1,965
+PubChem 6,912 571 1,162
+PCDes 7,474 1,051 2,136
+MoMu 7,474 1,051 4,554
 The statistics of datasets, including their the number of samples for training, validation and test sets are given in Table 5.
@@
 This innovative approach leverages transfer learning in the chemical domain, with a particular emphasis on addressing cross-domain tasks that involve both chemistry and natural language.
-Then, we will also include three baselines for unconditional generation to verify the effectiveness of our model.
+Then, we will also include seven baselines for unconditional generation to verify the effectiveness of our model.
+CharRNN (Segler et al., 2018): CharRNN models a distribution over the next token based on the previously generated tokens.
+We train this model by maximizing the log-likelihood of the training data, which is represented as SMILES strings.
+VAE (Kingma & Welling, 2013): Variational Autoencoder (VAE) consists of two neural networks: an encoder and a decoder.
+The encoder infers a mapping from high-dimensional data to a lower-dimensional space, while the decoder maps this lower-dimensional representation back to the original high-dimensional space.
+It optimizes reconstruction loss and regularization term in a form of Kullback-Leibler divergence.
+We use SMILES as input and output representations as discussed in (Polykovskiy et al., 2020).
+AAE (Makhzani et al., 2015): Adversarial Variational Autoencoder (AAE) replaces the Kullback-Leibler divergence in VAE with an adversarial objective.
+An auxiliary discriminator network is trained to distinguish between samples from a prior distribution and the model’s latent codes.
+Similar to VAE, we use SMILES as input and output representations as shown in (Polykovskiy et al., 2020).
+LatentGAN (Prykhodko et al., 2019): The Latent Vector Based Generative Adversarial Network (LatentGAN) integrates an autoencoder with a generative adversarial network.
+It first pretrains an autoencoder to map SMILES structures to latent vectors.
+Subsequently, a generative adversarial network is trained to generate latent vectors for the pre-trained decoder.
 BwR (Diamant et al., 2023): BwR aims to decrease the time/space complexity and output space of graph generative models, demonstrating notable performance improvements in generating high-quality outputs on biological and chemical datasets.
@@
 Utilizing these extracted principal subgraphs, it proceeds to generate molecules in two distinct phases.
+Note that we use the code for CharRNN, VAE, AAE, LatentGAN by the Molecular Sets (MOSES) Benchmark Polykovskiy et al. (2020).
 C.3 Setup and Hyper-parameter settings
@@
 For unconditional generation, we set $T$ as 10 or 5 for the sampling process.
-Appendix D More Visualization Results
-D.1 Property Conditional Molecule Generation
+Appendix D More Experimental Results
+D.1 Average Inference Time Comparison
+In this section, we compare the average inference times of 3M-Diffusion, MolT5, and ChemT5 by generating 1,000 samples, as shown in Figure 3.
+The average time to generate each molecule is calculated and presented.
+We observe that 3M-Diffusion achieves a speed 45 times faster than baseline methods, confirming the efficiency of our proposed model.
+Our proposed model not only generates molecules more quickly than previous approaches but also maintains a high level of semantic alignment with textual descriptions.
+D.2 Property Conditional Molecule Generation
 In this section, we provide more prompt examples and more experiment details for Property Conditional Molecule Generation.
-We generate 10 samples for each prompt and subsequently select the top 5 samples based on the desired property values, as shown in Figure 2,  6,  7 and  8.
+We generate 10 samples for each prompt and subsequently select the top 5 samples based on the desired property values, as shown in Figure 2,  5,  6 and  7.
 For instance, when using the prompt ”This molecule is not like a drug,” we select the top 5 molecules with the smallest QED values, as indicated.
@@
 Compared with the SOTA method MolT5-large, our generated results are more diverse and novel with maintained semantics in textual prompt.
-D.2 Qualitative Comparison
+D.3 Qualitative Comparison
 Figure 8: Qualitative comparisons to the MolT5-large in terms of generated molecules on CheBI-20.
@@
 We have the same observation that our model excels in generating molecules that not only closely resemble the reference molecule but also encompass some diverse and novel structures.
-D.3 Case Study
+D.4 Case Study
 In this section, we choose 20 more examples from the ChEBI-20 dataset to visually illustrate the structural similarity between the molecules generated by 3M-Diffusion and the ground truth molecules corresponding to the given text descriptions.
@@
 Figure 9: More molecules generated conditionally on input text by 3M-Diffusion on the ChEBI-20 dataset.
-D.4 Unconditional Molecule Generation
-We further present 50 molecules sampled from the prior distribution in Figure 11.
-As shown in Figure 11, our sampled molecules present rich variety and structural complexity.
-This demonstrates our model’s ability to generate diverse, novel, and realistic molecules.
+D.5 Unconditional Molecule Generation
+As shown in Figure 10, our sampled molecules present rich variety and structural complexity. This demonstrates our model’s ability to generate diverse, novel, and realistic molecules.
 Figure 10: 50 molecules sampled from the prior distribution $\mathcal{N}(\mathbf{0},\bm{I})$.